### v1

# Can one size fit all?: Measuring Failure in Multi-Document Summarization Domain Transfer

Abstract
Abstractive multi-document summarization (MDS) is the task of automatically summarizing information in multiple documents, from news articles to conversations with multiple speakers. The training approaches for current MDS models can be grouped into four approaches: end-to-end with special pre-training (“direct”), chunk-then-summarize, extract-then-summarize, and inference with GPT-style models. In this work, we evaluate MDS models across training approaches, domains, and dimensions (reference similarity, quality, and factuality), to analyze how and why models trained on one domain can fail to summarize documents from another (News, Science, and Conversation) in the zero-shot domain transfer setting. We define domain-transfer “failure” as a decrease in factuality, higher deviation from the target, and a general decrease in summary quality. In addition to exploring domain transfer for MDS models, we examine potential issues with applying popular summarization metrics out-of-the-box.
1 Introduction
Figure 1: Sample generations from SummN trained on Multi-XScience (SummN-Science) and PRIMERA trained on Multi-News (PRIMERA-News) on the ConvoSumm-Reddit dataset (Conversation). Relevance is measured automatically with UniEval.
Multi-document summarization (MDS) is an extended task of summarization, where the goal is to represent the key ideas from multiple documents in a single new document or paragraph. Current MDS systems (models) are trained and evaluated on data from a single domain, such as news, but there have been no robust studies on how models trained on these domains transfer, or perform, on different domains. While the overarching goal of MDS models is to condense information, models trained on one type of summary (e.g., news) often fail to transfer successfully to other domains (e.g., science) due to the unique nature of how summaries across different domains are written.
Recent MDS system surveys compare different models either on a single domain (e.g., conversation/dialogue (Zhang et al., 2021; Mullick et al., 2024)) or focus on MDS corpora biases (Wolhandler et al., 2022; Dey et al., 2020). While Zhang et al. (2021) found that retrieve-then-summarize models are best according to ROUGE, Mullick et al. (2024) did not find a clear pattern in model performance across a variety of metrics. Few works explore the zero-shot domain transfer ability of MDS systems, but Zhang et al. (2021) did find that pre-training on news datasets can transfer well to conversation. However, Fabbri et al. (2021a) a zero-shot model from news to conversation had lower performance than on conversation alone.
An example is in Figure 1, where two models, one trained on Multi-News (Fabbri et al., 2019) and the other on Multi-XScience (Lu et al., 2020) are evaluated on a sample from ConvoSumm-Reddit (Fabbri et al., 2021a). The difference in writing style is immediately apparent: the news-trained model draws out quotes, the science-trained model inserts citations, and both are completely different from the expected style of the Reddit conversation summary. The difference in style is also captured by similarity metrics ROUGE (Lin, 2004) and UniEval-Relevance (Zhong et al., 2022)–while the ROUGE scores, which capture word overlap, only differ by a few points, the semantic-based, trained Relevance scores vary widely.
While ConvoSumm-Reddit summaries are written to represent multiple points of view, other ways to represent online discourse is through an overall topic-based summary or detailed utterance-by-utterance summary. All types would “condense” the same conversation but at varying levels of granularity, style, and word choice. Similarly, summaries for other domains also vary in incorporated information and style.
Figure 1 highlights the three axes of domain transfer for MDS systems: 1) domain of the training data, 2) type of model, and 3) evaluation metrics. While there are popular benchmarks for summarization, including the notable SummEval (Fabbri et al., 2021b), few are for multi-document summarization, and most datasets focus on the news domain. We choose three high-quality human-written datasets to represent News (Multi-News (Fabbri et al., 2019)), Science (Multi-XScience (Lu et al., 2020)), and Conversation (ConvoSumm-Reddit (Fabbri et al., 2021a)) domains.
Regarding the type of model, while there are many models for summarization, the types of models for abstract multi-document summarization are end-to-end with special pre-training (“direct”), chunk-then-summarize, extract-then-summarize, and direct with large language model (LLM) decoders. There are many models within each type (Mullick et al., 2024) and for greater in-depth analyses, we choose one, strong representative model of each kind. We focus on neural, transformer models and omit extractive comparisons due to previous work demonstrating the stronger performance of neural models (Fabbri et al., 2021b).
Similarly, while there are many evaluation metrics for generated text, which each have their strengths and weaknesses (He et al., 2023; von Däniken et al., 2022; Fabbri et al., 2021b), we group commonly used metrics into three categories: 1) similarity to summary reference (ground-truth), 2) overall quality (e.g., fluency), and 3) groundedness in the source documents. We find that: [^1] All code and models will be released upon publication.
  • Smaller language models (BART-size) hold more stylistic bias with their training data;
  • Within News, Science, and Conversation, there is no clear “best” domain for general transfer;
  • Lexical and semantic similarity metrics both fail to capture summary quality;
  • Current non-reference-based metrics do not scale well to the multi-document setting;
  • Trained evaluators have bias and penalize out-of-domain datasets;
Figure 2: Summary of all the domains (datasets), multi-document summarization (MDS) model types, and evaluation metrics analyzed in this work.
2 Domain Transfer of Multi-Document Summarization
While there is in-depth work on analyzing summarization and MDS systems and corpora (Zhang et al., 2021; Mullick et al., 2024; Fabbri et al., 2021b; Dey et al., 2020), there is no systematic evaluation of how each model type performs when applied out-of-domain and how the automatic evaluation metrics can over- or under-represent model performance. We draw from existing models and tools to explore the intricacies of domain transfer with MDS models and their evaluation methods. The main dimensions of our analyses are domain (i.e., dataset), model type, and evaluation metrics. See Figure 2 for summary.
2.1 Summarization Dataset Domains
News is a common domain for summarization datasets (Hermann et al., 2015; Fabbri et al., 2019; Jiang and Dreyer, 2024; Choi et al., 2024; Gholipour Ghalandari et al., 2020; Litkowski and Road, 2004) There are datasets for other domains like scientific articles (Lu et al., 2020) and conversation (Fabbri et al., 2021a; Gliwa et al., 2019; Chen et al., 2021). We select a single, high-quality multi-document dataset with human-written summaries to represent each domain. All documents and summaries are in English.
Conversation.
We use the Reddit subset of the released data from ConvoSumm (Fabbri et al., 2021a) (ConvoSumm-Reddit). Since the summaries were written by paid crowdsource workers, this dataset is significantly smaller than the news and science datasets, with only 500 entries.
News.
Multi-News (Fabbri et al., 2019) is a news-based MDS dataset featuring 56,216 human-written summaries from Newser.222newser.com
Science.
Representing the Science domain, the entries in Multi-XScience pair a related works section of an academic paper with 1) its abstract and 2) the abstracts from cited work. The 40,528 papers were collected from arXiv and the Microsoft Academic Graph (MAG) (Lu et al., 2020). Since the paper dump was from pre-2020, we assume all writing on arXiv is human-generated.
2.2 Multi-Document Summarization Model Types
Current high-performing fine-tuned abstractive summarization methods can be grouped into four approaches: summary-specific pre-training (Zhang et al., 2020a; Xiao et al., 2022a; Peper et al., 2024), incremental summarization (“chunk-then-summarize”) (Zhang et al., 2022), extracting claims or viewpoints (“extract-then-summarize”) (Fabbri et al., 2021a; Ernst et al., 2022; Ouyang et al., 2023), and fine-tuning or few-shot summarization with large pre-trained models (Mullick et al., 2024). These groupings have been used in prior work (Mullick et al., 2024).
Modified Pretraining.
Prior work has shown that fine-tuning a language model specifically pre-trained with summary-focused objectives improves downstream performance over non-specialized models (Xiao et al., 2022b; Zhang et al., 2020a; Peper et al., 2024). Zhang et al. (2020a) first introduced this concept of summary-specific pre-training, and it was later augmented to the multi-document case (Xiao et al., 2022b; Peper et al., 2024). We select PRIMERA to represent this group. PRIMERA is Longformer-Encoder-Decoder model initialized from BART-Large (Beltagy et al., 2020) and pre-trained with an Entity Pyramid objective on a large news corpora (Xiao et al., 2022b). While PELMS (Peper et al., 2024) adjusts the pre-training strategy to improve upon PRIMERA, we omit comparison to this model because it also uses a data filtering step, which overlaps with the extract-then-summarize approach.
Chunk-then-summarize.
Another approach to MDS is incremental summarization of documents, as done with SummN (Zhang et al., 2022). SummN first segments the source documents and target summary and creates mappings between the source and target segments (via ROUGE max similarity), which and trains a separate BART-Large-CNN model (Zhang et al., 2021) each “coarse” stages before the final “fine-grained” stage which produces the end summary.
Extract-then-summarize.
A different approach to the long-context problem of MDS systems is to first extract the salient portions of the documents before summarizing them. The main models in this category are COMPO (Ouyang et al., 2023), ProCluster (Ernst et al., 2022), and ConvoSumm (Fabbri et al., 2021a),. COMPO is specifically tailored toward conversational summarization and provides summaries based on extracted sub-components (topics and actions) within the dialogue. ProCluster extracts “propositions” in documents, clusters them, and then summarizes the clusters. ConvoSumm first filters sentences from documents to only the “claims” and then summarizes the claims. A variation of the extract-then-summarize pipeline is attribute-then-generate (Slobodkin et al., 2024) which requires multiple steps of content selection, sentence planning, and sentence-by-sentence generation. We chose ConvoSumm over other extract-then-summarize models due to its applicability to many domains and non-overlap with different approaches with clustering or sentence-planning steps.
Pre-trained LLM.
While the above models are strong MDS systems, they are all the size of BART-Large (336M params.). A growing trend across natural language processing research has been to fine-tune or adapt large, pre-trained decoder language models (LLMs) through in-context learning. We select Llama-3.1-8B-Instruct to represent this approach due to its strong performance on a variety of tasks (Dubey et al., 2024).
2.3 Summarization Evaluation
Most MDS papers and their surveys use off-the-shelf automatic evaluation metrics without discussing drawbacks. Summarization metrics usually measure one or more of the following attributes: similarity to a ground-truth summary, overall quality, and factuality. While many MDS systems are evaluated with manual annotations, we focus on automatic evaluation in this work.
Similarity to Ground Truth.
Most text evaluation metrics compare the model output to a known, quality summary or ground truth (i.e., reference). The most common metric for evaluating summary quality is ROUGE (Lin, 2004) but is better viewed as a similarity metric since it is the token recall of the generated summary based on the reference (i.e., ground-truth). The complement of ROUGE is BLEU (Papineni et al., 2002), which measures precision instead of recall of generated summaries.
Other metrics in the spirit of ROUGE/BLEU are semantic-based instead of lexical, which do not penalize generations with similar meanings but different word choices. The most popular version is BERTScore (Zhang et al., 2020b), which maps tokens from the generation to the reference before calculating overlap. An alternative is MoverScore, which “optimized” version of BERTScore (Zhang et al., 2020b) which uses soft token alignments from the hypothesis to the reference instead of a hard 1:1 mapping, allowing for a greater reward for semantically similar generated text.
Other similarity metrics are sub-“dimensions” of a trained multi-dimensional evaluator, like UniEval. UniEval is a T5-Large model fine-tuned on synthetically augmented datasets for positive and negative examples of different “dimensions” of summary attributes (fluency, relevance, coherence, consistency) (Zhong et al., 2022). The UniEval-Relevance dimension is a similarity metric that compares the reference and generation as a whole instead of per token.
Overall Quality.
Not all desired qualities of a summary depend on a reference, some are intrinsic or rely on the source documents (e.g., fluency). UniEval measures both fluency and coherence. UniEval-Fluency is based solely on the generation, and UniEval-Coherence measures the coherence of the generation with respect to the source.
Factuality.
A growing concern with generated text is the possibility of hallucinations and non-faithfulness to the source material (Min et al., 2023). While objective factuality is important for the safety of LLMs (e.g., knowing that boiling water is hot), for MDS we typically assume that the source documents are factually correct and that a summary is only factual if its claims match the claims from the source. We measure summary factuality with UniEval-Consistency, a single-dimensional evaluator trained to measure the factual consistency of generated text to the source (as opposed to continuously trained as with UniEval-Fluency and UniEval-Coherence). A popular framework for factuality analysis is decompose-then-verify (Min et al., 2023; Jiang et al., 2024), and we leave this analysis to future work.
We include ROUGE, MoverScore, and multiple dimensions of UniEval in our analysis as representations for similarity to ground truth, factuality, and overall quality metrics. While BARTScore (Yuan et al., 2021) measures many of the same dimensions as UniEval, we chose UniEval because it is better correlated with human evaluation (Zhong et al., 2022) and the potential bias of BARTScore to model generations, since most of the models are BART-based (“self bias”) (He et al., 2023). Another evaluation for summaries is source coverage, or how much of the source documents are represented in the summaries (Wolhandler et al., 2022). We leave this evaluation to future work.
3 Experimental Setup
Our goal is to explore the intricacies of domain transfer with multi-document summarization models and their evaluation methods. The main dimensions of our analyses are different datasets (i.e., domains), models, and evaluation metrics (Figure 2). We chose 3 dataset domains (News, Science, Conversation), 4 model types (end-to-end/direct), chunk-then-summarize, extract-then-summarize, and pre-trained LLMs), and 7 metrics for a total of 84 points of analysis (Section 2).
3.1 Model Training
The hyperparameters for each model and dataset combination are in Appendix B. Each model was trained on each dataset, and we refer to them as ModelName-Domain, e.g., ConvoSumm fine-tuned on Multi-News is ConvoSumm-News.
PRIMERA.
As input, the source documents are delineated with a special <doc-sep> token, and the model’s global attention is set on those tokens. To handle examples with source documents that are too long for the context window, we follow the author’s code and evenly truncate the documents to fit. We used the authors’ released models of PRIMERA fine-tuned on Multi-XScience and MultiNews and trained PRIMERA-ConvoSumm-Reddit with the released author code.
SummN.
We use a 3 course-stage process for all datasets. A separate model is trained for each coarse stage, leading to $N+1$, or $4$ fine-tuned BART models for each training dataset. An important note is due to the setup of summarizing chunks, there are no too-long context issues as with direct models like PRIMERA and all source documents are considered by the model.
ConvoSumm.
There are multiple versions of ConvoSumm but we select ConvoSumm-argument-filtered (“arg-filtered”) due to its comparable performance to the more complex argument-graph version. The argument filtration model is a BERT model fine-tuned on annotated data from the Change My View community on Reddit (i.e., conversation, AMPERSAND (Chakrabarty et al., 2019)) and annotated persuasive essays on a variety of topics (Stab and Gurevych, 2017)). A BART-Large model with an expanded context window (2048) is fine-tuned on the extracted claims.
Llama 3.1.
Unlike the other models, we do not fine-tune Llama 3.1 8B-Instruct. [^3] https://huggingface.co/meta-llama/Llama-3.1-8B-Instruct To compare to the domain-transfer experiments, we provide an example from the training set to the model, i.e., for News-Convo, an example from Multi-News is provided to Llama and then the model is prompted to complete an example from ConvoSumm. We also include a zero-shot version which does not provide a complete example to Llama.
3.2 Evaluation Details
ROUGE scores have been known to vary depending on the implementation used. We use the HuggingFace wrapper for the Google Research implementation. [^4] https://huggingface.co/spaces/evaluate-metric/rouge We report ROUGE-1 (R1) and ROUGE-L-Sum (RL-Sum) to follow prior work. For the other evaluation metrics, MoverScore and multiple dimensions from UniEval, we use the trained models as they are without modification.
Factuality Ground-Truth Similarity Quality
Cons. (UE) Rel. (UE) R1 RLSum MoverScore Fluency (UE) Coh. (UE)
Conversation 0.68 0.75 0.30 0.20 0.07 0.91 0.72
Avg. $\Delta$ +18% -13% -9% -19% -31% -1% +4%
Ref. 0.71 0.99 1.00 1.00 1.00 0.90 0.73
News 0.72 0.78 0.43 0.22 0.20 0.91 0.79
Avg. $\Delta$7% +2% -51% -41% -92% -2% +1%
Ref. 0.72 0.99 1.00 1.00 1.00 0.92 0.80
Science 0.78 0.63 0.24 0.15 0.03 0.78 0.64
Avg. $\Delta$7% -12% +6% +1% -7% -1% +4%
Ref. 0.55 0.99 1.00 1.00 1.00 0.66 0.43
Table 1: Multi-document summarization (MDS) model performance and relative performance aggregated by domain when trained and tested on the same domain (top) and different domains ($\Delta$), respectively. Results are aggregated across all the test splits of each dataset. A positive $\Delta$ shows a performance gain and a negative $\Delta$ is a performance loss as compared to the model’s scores on the in-domain data. A relative change of over $1$ is possible with MoverScore because of its $[-1,1]$ range. Detailed results are in Table 7. “Ref” refers to the results when running the metrics on the ground-truth data.
4 Results
We analyze the results in the framework of our three axes: dataset domain, model type, and evaluation metric. For ease of analysis, we show the relative performance of each model when evaluated on a new domain in Tables 2 and 1. The relative performance is the percentage gain or loss of a model’s performance when compared to the performance on the in-domain data. The zero-shot baseline with Llama 3.1 is compared to the in-context Llama 3.1 performance for each dataset (i.e., no examples instead of one example). See Appendix C for the detailed per-model and dataset results.
4.1 Domain Generalizability
When aggregated by dataset over the different model types, clear patterns emerge from Table 1.
Science summaries have the lowest general quality.
The quality is lowest in the Science domain compared to Conversation and News according to Fluency and Coherence metrics, even before evaluating Science-trained MDS models on other domains, $0.78$ and $0.64$, respectively. These scores are $10+$ points lower than those of News ($0.91$/$0.79$) and Conversation ($0.91$/$0.72$).
Fluency is affected the least by domain transfer.
The fluency scores are the most stable, i.e., have the least change in performance (Avg $\Delta$) compared to the other metrics. While on average there is a slight decline in fluency, as indicated by the negative change, the decline is negligible in the $(-0.02,-0.01)\%$ range. Fluency does not appear to be an issue with these models, mostly likely due their size (BART-Large and Llama 8B).
Factuality Ground-Truth Similarity Quality
Cons. (UE) Rel. (UE) R1 RLSum MoverScore Fluency (UE) Coh. (UE)
ConvoSumm 0.65 0.62 0.32 0.18 0.08 0.83 0.57
Avg. $\Delta$2% -11% -28% -26% -152% +3% -4%
PRIMERA 0.78 0.74 0.35 0.20 0.12 0.89 0.78
Avg. $\Delta$ +3% -10% -19% -24% -16% -14% -1%
SummN 0.66 0.60 0.33 0.19 0.08 0.81 0.56
Avg. $\Delta$ +4% -9% -32% -33% -31% +7% +18%
Llama 3.1 0.82 0.91 0.29 0.18 0.11 0.94 0.94
Avg. $\Delta$ +1% -1% +7% +4% +25% -0.00% +0.00
Llama 3.1 (0-shot) +1% -4% -5% +0% -7% -1% -1%
Table 2: Multi-document summarization (MDS) model performance and relative performance aggregated by model when trained and tested on the same domain (top) and different domains ($\Delta$), respectively. Results are aggregated across all the test splits of each dataset. A positive $\Delta$ shows a performance gain and a negative $\Delta$ is a performance loss as compared to the model’s scores on the in-domain data. A relative change of over $1$ is possible with MoverScore because of its $[-1,1]$ range. LLama zero-shot is compared to Llama with in-domain examples. Detailed results are in Tables 5 and 6.
There is no clear general-purpose domain for transferring.
Each domain has different benefits and cons to train on for transferring, as indicated by the varied relative performance scores ($\Delta$). Transferring from Conversation to News and Science can lead to greatly increased factuality, as measured by UniEval-Consistency ($+18\%$), but also a decline in Ground-Truth Similarity ($-31\%$ to $-9\%$). Transferring from News appears to be the worst according to the similarity metrics, over $-92\%$ according to MoverScore. The Science-trained models appear to be the most neutral overall, ranging from $-12\%$ to $+6\%$ across all metrics, but claiming that it is the best domain for general MDS model training is difficult because the model scores started lower than the other domains to begin with.
News-trained models have the worst similarity when transferring to non-News data.
There is a steep decline in performance when News-trained models are evaluated on Conversation and Science as indicated by ROUGE ($-51\%$ R1, $-41\%$ RL-Sum) and MoverScore ($-92\%$). However, the Relevance score is high ($+2\%$). The discrepancy in word-mapping metrics (lexical with ROUGE and semantic with MoverScore) and overall similarity (Relevance) show that a summary can still incorporate information from the ground-truth without having an overlap in words.
Conversation and News data have the most unique styles.
The ConvoSumm-Reddit example from Figure 1 shows the stark difference in style between the types of summaries in ConvoSumm-Reddit and those from Multi-News. This difference in style explains the performance loss from Conversation and News domains, as the summary wording would be completely different and penalized by the ground-truth similarity metrics ($-31$ to $-9\%$ and $-92$ to $+2\%$, respectively).
4.2 MDS Model Transfer
Grouping the domain transfer results by model type across all dataset domains reveals even more performance trends (Table 2).
Pre-trained LLM outperforms all fine-tuned smaller models.
For all metrics besides similarity, Llama 3.1 outperforms the other models concerning factuality and quality. This trend holds during domain transfer, with a performance difference ranging from $-1$ to $+25\%$ across all metrics. The second-best performing model across all metrics, including accuracy, is PRIMERA, indicating an advantage of directly fine-tuning a long-context model (4K context window).
Zero-shot performance is the same as in-context learning with Llama 3.1.
Surprisingly, Llama has a similar performance at generating summaries with (in-context learning, ICL) or without being shown an example (zero-shot). This indicates a strong bias to its pre-training versus “learning” from the example. These results could differ with the addition of more examples (at the expense of a smaller context window available for the source documents) or a larger model.
ConvoSumm claim extraction is ineffective on Science.
The first step in the extract-then-summarize model ConvoSumm is to filter out sentences that are not identified as claims by a small, fine-tuned model. This step removes $35\%$ of sentences on average from each ConvoSumm-Reddit example, $51\%$ from Multi-News, but less than $1\%$ from Multi-XScience. This is intuitive since there would be more filler sentences in conversations and news articles, but less so in scientific publications.
Chunk-then-summarize is the only approach guaranteed to fit all source documents.
With all models, including Llama 3.1 with a 131K context window, there were issues fitting all source documents for every example comfortably in the context windows. As mentioned in Section 3, PRIMERA “handles” this issue by evenly dividing the context window among the source documents and truncating the ends. The assumption that the beginning of documents is more important than the end does not hold for all domains and has only been identified in News (Dey et al., 2020). While ConvoSumm was able to fit all source documents from ConvoSumm-Reddit, it did have to truncate documents from 25 samples in Multi-XScience and 699 samples from MultiNews. PRIMERA, with a context window double the size of ConvoSumm (2K vs 4K), was able to fit all Conversation and Science samples, but truncated 835 Multi-News examples. Llama only had issues with 3 samples from Multi-News.
4.3 Evaluation of Evaluation
Word-based similarity scores are low across all models.
For all models, the generated summaries are not very similar to the ground-truth reference according to the lexical and semantic similarity metrics ROUGE and MoverScore. The model with the highest accuracy ($0.20$ ROUGE_L and $0.12$ MoverScore) was PRIMERA, higher than the much larger LM, Llama-8B. For lexical-based metrics like ROUGE, the low similarity can be explained through the model’s use of synonyms, which would not be captured through token-based recall. However, the low MoverScores can be explained by (1) the penalization of generations longer or more detailed than the reference and (2) the lack of a normalizing constant as provided by BERTScore (Zhang et al., 2020b). [^5] Zhang et al. (2020b) use a learned constant to normalize the scores to $[0,1]$ for easier readability. Still, low scores across both accuracy metrics indicate that the generated summaries differ from their human-written counterparts. Despite overall low accuracy scores, all models have the highest ROUGE/MoverScore performance on Multi-News (News), indicating the summaries might not be as abstract as the other datasets.
Surprisingly, the Relevance score which compares the reference and generation as a whole instead of token mappings, shows higher similarity, especially for Llama ($0.91$ average).
Accuracy is affected the most.
The reference-based metric, accuracy, shows the most extreme gains and losses compared to the other metrics. This is intuitive because a model’s summary can still be good but not be written in the same manner, or style, as the reference, which can happen if a model is biased towards the writing style of its training data. The zero-shot baseline, which is only prompted and not given any styles, shows the least variance in accuracy compared to the other models.
All models are fluent.
Unsurprisingly, all models have high fluency scores ($0.80+$) since all models are of size BART-Large and above. While fluency is still an important attribute to check, it might not be as important as it once was.
Non-reference metrics do not scale well to the multi-document setting.
Evaluating for attributes that depend on the source documents (Consistency and Coherence) is most affected by the long-document nature of multi-document summarization. The options are either to 1) truncate the documents, which can greatly alter the scores depending on which documents are included, or 2) calculate the metric for every source document and aggregate the scores. More work is needed to determine whether truncating the source documents for UniEval over- or under-estimates model performance. Another option is to turn to large-context models like GPT-4 to be a “judge” for specific criteria (Kim et al., 2024; Vu et al., 2024). This use of LLMs, which typically can fit 10K+ tokens in their context windows, is a potential solution to this scaling issue. The reference-based metrics, i.e., similarity metrics that only compare the ground-truth and generated summaries, are not affected by this issue.
Trained metrics can penalize out-of-training data.
Just as summarization models can have lower performance on out-of-domain data, so can trained evaluation models like UniEval. As noted before, the fluency and coherence scores on Science are $0.10$ points lower than on News and Conversation. This could be an issue with the summarization model training, but the fluency and coherence scores are also lower on the ground-truth summaries, $0.30+$ points lower (“Ref.” in Table 1), which should not be the case as all summaries are human-written and should have high scores. An explanation is that UniEval was trained on a synthetically augmented version of news data, CNN/DailyMail, which is very different from Scientific text.
5 Discussion and Conclusion
In most work that surveys or proposes multi-document summarization (MDS) systems, there is little overlap between the analysis of model performance and the shortcomings of popularly used summarization evaluation metrics. We bridged this gap through experiments and analysis of not only the zero-shot domain transfer ability of MDS systems but also their evaluation. We analyzed three components of a multi-document summarization (MDS) pipeline for domain transfer: 1) the domain of the training data, 2) the model type, and 3) the common evaluation metrics.
We found important facets to consider in each component. For domain transfer, the training and test domains are important for model performance, regardless of the model type. While we did not find a clear trend between Conversation, News, and Science, it was apparent that the transfer was highly affected by the style of the summaries, albeit more than the domain itself. Summaries from different domains but written in the same style (e.g., high-level vs low-level) could transfer better with higher similarity to the reference.
For different approaches to MDS, a model that didn’t hold a strong bias towards the style of its training (or example) data is ideal, as is the case with a large pre-trained decoder model, like Llama 3.1. If the domain styles differ greatly, a zero- or few-shot approach with less bias is ideal.
Notable issues we uncovered with evaluation metrics that prevented the fair evaluation of summaries were domain-specific penalizations, mostly for Science, questions with scaling to the MDS case, and the inability of ground-truth similarity metrics to adjust for style change (ROUGE and MoverScore). When evaluating an MDS model on a new domain, word-similarity metrics judge style more than content similarity and trained evaluators could over- or under-judge model performance due to limited information in the context window.
Limitations
Despite the rise in popularity of large, pre-trained decoder models, most fine-tuned multi-document summarization models are on the scale of a few hundred million parameters, e.g., BART-Large. We did not analyze the effect of model size or scale concerning summarization performance or domain transfer abilities.
Another limitation is our focus on English, as all of the datasets and learned metrics (i.e., UniEval), were in English or trained on English data.
Ethical Considerations
We did not evaluate the model output and training datasets for absolute factuality (i.e., correctness with respect to the world) or harmful language. We leave these tasks to future work or the incorporation of an additional step for model safety.
Appendix A Dataset Details
We include three MDS datasets: Multi-News (Fabbri et al., 2019), Multi-XScience (Lu et al., 2020)), and ConvoSumm-Reddit (Fabbri et al., 2021a). We keep the original train, validation, and test splits as the original papers for all datasets. See Table 3 for an example from each dataset and Table 4 for summary statistics.
Multi-News.
For Multi-News we used the “cleaned” version provided by the authors [^6] https://github.com/Alex-Fabbri/Multi-News/issues/36 with 44964/5621/5621 examples for train, validation, and test sets, respectively. We keep the NEWLINE_CHAR symbol in the documents.
Multi-XScience.
The Multi-XScience dataset had many examples with missing source documents. In addition to skipping those with no source documents, we also removed non “multi” examples which had less than two source documents, resulting in 11523/1937/1901 examples for train, validation, and test sets, respectively. We skipped 18846/3129/3192 examples from train/val/test sets with less than two source documents (i.e., abstracts) and 20023/3383/3403 examples from train/val/test with completely missing source documents. We also replace all document-specific citations cite_5 with a more generalizable cite.
ConvoSumm.
While ConvoSumm features multiple conversation datasets, we focus on Reddit. Similar to the original work, we flatten comment threads as individual source documents, resulting in 201/50/250 examples for train, validation, and test sets, respectively. We keep the |NEWLINE| symbols in the documents.
Appendix B Training Details
Unless specified otherwise, each model was trained on a single Nvidia A100 GPU with 80GB of memory. The training hyperparameters for PRIMERA, SummN, and ConvoSumm-arg-filtered are in Tables 10, 9 and 11, respectively. Below are the prompts and details for running Llama for the zero-shot and in-context 1-shot comparisons.
PRIMERA
The authors’ released models were from HuggingFace fine-tuned on Multi-XScience (https://huggingface.co/allenai/PRIMERA-multixscience) and MultiNews (https://huggingface.co/allenai/PRIMERA-multinews). The authors’ code is on GitHub (https://github.com/allenai/PRIMER).
Llama 3.1 8B-Instruct.
Due to the large size of Llama 8B, we hosted it as a service with vLLM on 4 Nvidia A100 80GB GPUs (Kwon et al., 2023). [^7] https://docs.vllm.ai We followed model examples on how to interact for the summarization task (Table 12).
Appendix C Raw Results
The domain-transfer results in Section 4.2 show the summarized relative performance of the models. The detailed performance across domains and models are in Tables 7 and 6, respectively. The raw, un-normalized numbers are shown in Table 8.
Dataset Source Documents Summary
Multi-News Whether a sign of a good read; or a comment on the ’pulp’ nature of some genres of fiction, the Oxfam second-hand book charts have remained in The Da Vinci Code author’s favour for the past four years. NEWLINE_CHAR NEWLINE_CHAR Dan Brown has topped Oxfam’s ’most donated’ list again, his fourth consecutive year …
A charity shop is urging people to stop donating The Da Vinci Code after becoming overwhelmed with copies. NEWLINE_CHAR NEWLINE_CHAR The Oxfam shop in Swansea has been receiving an average of one copy of the Dan Brown novel a week for months, leaving them with little room for any other books. NEWLINE_CHAR NEWLINE_CHAR Staff who are struggling to sell copies of the book have put a note up in the store saying they would rather donors hand in their vinyl instead.
– The Da Vinci Code has sold so many copies—that would be at least 80 million—that it’s bound to turn up in book donation piles. But at one charity shop in the UK, it’s been donated so heavily that the shop has posted a sign propped up on a tower of Da Vinci Code copies that reads: Ÿou could give us another Da Vinci Code… but we would rather have your vinyl!T̈he manager of the Oxfam shop in Swansea tells the Telegraph that people are laughing and taking pictures of the sizable display: Ï would say that we get one copy of the book every day.Ḧe says people buy them öccasionally,b̈ut with vinyl sales up 25% in the past year, they’d rather take records. Dan Brown’s book isn’t the only one that shops like Oxfam struggle to re-sell. Last year, Oxfam was hit with a large and steady supply of Fifty Shades of Grey, and it similarly begged donors: P̈lease—no more.B̈ut Brown has a particular kind of staying power. The Da Vinci Code was published in 2003, and within six years Brown had booted John Grisham from the No. 1 slot on the list of writers whose books were most often donated to Oxfam’s 700 shops, reported the Guardian at the time. The Independent in 2012 reported Brown’s best-seller was the most-donated book for the fourth year running. (See why Dan Brown took heat from the Philippines.)
Multi-XScience Despite the huge amount of recent research efforts on entity resolution (matching) there has not yet been a comparative evaluation on the relative effectiveness and efficiency of alternate approaches. We therefore present such an evaluation of existing implementations on challenging real-world match tasks. We consider approaches both with and without using machine learning to find suitable parameterization and combination of similarity functions. …
This tutorial provides a comprehensive and cohesive overview of the key research results in the area of record linkage methodologies and algorithms for identifying approximate duplicate records, and available tools for this purpose. It encompasses techniques introduced in several communities including databases, information retrieval, statistics and machine learning. It aims to identify similarities and differences across the techniques as well as their merits and limitations.
For accurate entity linking and deduplication, early works on the subject examined many methods such as cosine similarity match, distance based match, TF IDF, and Soundex. The well known similarity measures for entity linking are summarized and reviewed in cite ; and the work in cite presents a comparative evaluation of some existing works.
ConvoSumm Title: Hand from tonight. Subreddit: poker Playing at my local pub league. Villian is loose aggressive. Sorry if this is bad formatting just quickly whipped it up for your opinions. |NEWLINE| Button: ( 30000) |NEWLINE| SB: Villian: (50000) …
Shove pre but I imagine this is the sort of game where you get told off for doing that
A lot of commenters debate the semantics and word choice of another comment. Some commenters say that the poster should shove even though other players will not like that move very much. Other commenters discuss the differences of some poker terminology.
Table 3: Examples from each of the three datasets: Multi-News, Multi-XScience, and ConvoSumm-Reddit.
# Examples # Source Docs. Summary Length Avg. Doc. Length
ConvoSumm (Reddit) Test 250 14.42 (5.55) 57.78 (12.04) 53.02 (26.60)
Train 201 14.98 (6.28) 58.32 (12.95) 49.46 (26.38)
Val 50 16.74 (7.21) 59.18 (14.99) 52.88 (27.29)
MultiNews Test 5621 2.73 (1.08) 217.00 (67.20) 982.98 (1779.55)
Train 44964 2.76 (1.12) 218.27 (68.04) 982.95 (1851.20)
Val 5621 2.70 (1.04) 216.72 (67.67) 980.06 (1163.94)
MultiXScience Test 1901 3.55 (1.95) 106.35 (41.60) 195.32 (48.01)
Train 11523 3.58 (2.05) 106.82 (42.14) 197.03 (47.91)
Val 1937 3.55 (1.99) 104.03 (41.38) 196.98 (48.84)
Table 4: Statistics for each dataset used in this work: Multi-News, Multi-XScience, and ConvoSumm-Reddit. The number of tokens in the (ground truth) summary and documents are measured with the BART tokenizer. Since we remove examples with missing information, our numbers might not match those in the original dataset papers (Appendix A).
Domain Factuality Accuracy Quality
Train/Test Cons. (UE) Rel. (UE) RL-Sum MoverScore Fluency (UE) Coh. (UE)
ConvoSumm Convo 0.64 0.70 0.21 0.06 0.91 0.60
News 0.64 0.62 0.19 0.17 0.89 0.66
Science 0.67 0.55 0.15 0.01 0.69 0.46
All 0.65 0.62 0.18 0.08 0.83 0.57
PRIMERA Convo 0.65 0.73 0.21 0.09 0.92 0.71
News 0.74 0.82 0.25 0.21 0.93 0.78
Science 0.96 0.68 0.14 0.05 0.83 0.86
All 0.78 0.74 0.20 0.12 0.89 0.78
SummN Convo 0.65 0.63 0.21 0.07 0.89 0.62
News 0.69 0.78 0.21 0.20 0.89 0.80
Science 0.65 0.39 0.15 -0.02 0.66 0.27
All 0.66 0.60 0.19 0.08 0.81 0.56
Llama 3.1 Convo 0.80 0.92 0.14 0.07 0.93 0.94
News 0.82 0.90 0.19 0.19 0.94 0.93
Science 0.85 0.90 0.12 0.08 0.94 0.95
All 0.83 0.91 0.13 0.08 0.93 0.95
Table 5: Multi-document summarization (MDS) model performance when trained and tested on the same domain. The results for “All” are averaged across all domains for each model. Results are on the test splits of each dataset. Metrics followed by (UE) are from UniEval. For all metrics a greater score is better.
Domain Factuality Accuracy Quality
Train Test Fact. (UE) Rel. (UE) R1 RLSum MoverScore Fluency (UE) Coh. (UE)
ConvoSumm Convo News 0.05 -0.41 -0.37 -0.44 -1.67 0.00 -0.25
Convo Science 0.06 -0.09 -0.25 -0.33 -0.72 -0.01 0.02
News Convo -0.11 0.30 -0.46 -0.31 -1.12 0.02 0.19
News Science -0.07 0.10 -0.41 -0.28 -0.90 0.05 0.02
Science Convo -0.07 -0.30 -0.16 -0.10 -2.79 0.03 -0.13
Science News -0.01 -0.28 -0.01 -0.08 -1.90 0.07 -0.07
All 0.01 -0.02 -0.11 -0.28 -0.26 -1.52 0.03 -0.04
PRIMERA Convo News 0.20 -0.31 -0.05 -0.22 -0.22 -0.09 -0.08
Convo Science 0.48 -0.07 -0.24 -0.36 -0.47 -0.10 0.20
News Convo -0.22 0.00 -0.61 -0.57 -1.03 -0.07 0.00
News Science 0.29 -0.17 -0.48 -0.45 -0.78 -0.11 0.09
Science Convo -0.47 -0.12 -0.20 -0.13 -0.74 -0.27 -0.25
Science News -0.11 0.04 0.43 0.30 2.25 -0.21 -0.03
All 0.03 -0.10 -0.19 -0.24 -0.16 -0.14 -0.01
SummN Convo News 0.31 -0.08 -0.22 -0.30 -0.67 0.04 0.14
Convo Science 0.26 -0.02 -0.31 -0.35 -0.79 0.04 0.31
News Convo -0.32 0.05 -0.69 -0.60 -1.22 -0.00 -0.00
News Science -0.17 -0.21 -0.52 -0.41 -1.00 -0.01 -0.25
Science Convo -0.00 -0.34 -0.59 -0.51 5.56 0.05 0.14
Science News 0.17 0.03 0.37 0.17 -3.74 0.26 0.72
All -0.06 0.04 -0.09 -0.32 -0.33 -0.31 0.07 0.18
Llama 3.1 Convo News 0.04 -0.04 0.69 0.42 1.74 0.01 -0.01
Convo Science 0.07 -0.03 0.05 0.03 0.29 -0.00 0.03
News Convo -0.02 0.04 -0.48 -0.36 -0.72 -0.01 0.03
News Science 0.05 -0.00 -0.42 -0.30 -0.56 -0.01 0.04
Science Convo -0.06 0.02 -0.10 -0.05 -0.34 -0.01 -0.02
Science News -0.03 -0.05 0.70 0.49 1.11 0.01 -0.05
All 0.01 -0.01 0.07 0.04 0.25 -0.00 0.00
Llama 3.1
(0-shot)
zero-shot Convo -0.01 -0.03 -0.16 -0.07 -0.23 -0.01 -0.02
zero-shot News 0.02 -0.07 -0.01 0.06 -0.01 -0.00 -0.03
zero-shot Science 0.01 -0.01 0.01 0.02 0.01 -0.01 0.01
All 0.01 -0.04 -0.05 0.00 -0.07 -0.01 -0.01
Table 6: Multi-document summarization (MDS) model relative performance when trained and tested on different domains. Results are on the test splits of each dataset. A positive number shows a performance gain and a negative number is a performance loss as compared to the model’s scores on the in-domain data. A relative change of over $1$ is possible with MoverScore because of its $[-1,1]$ range.
Domain Factuality Accuracy Quality
Train Test Fact. (UE) Rel. (UE) R1 RLSum MoverScore Fluency (UE) Coh. (UE)
Convo News 0.15 -0.21 0.01 -0.14 -0.20 -0.01 -0.05
Science 0.22 -0.05 -0.19 -0.25 -0.42 -0.02 0.14
All 0.18 -0.13 -0.09 -0.19 -0.31 -0.01 0.04
News Convo -0.17 0.10 -0.56 -0.46 -1.02 -0.02 0.05
Science 0.03 -0.07 -0.46 -0.36 -0.81 -0.02 -0.03
All -0.07 0.02 -0.51 -0.41 -0.92 -0.02 0.01
Science Convo -0.15 -0.18 -0.26 -0.20 0.42 -0.05 -0.06
News 0.01 -0.06 0.37 0.22 -0.57 0.03 0.14
All -0.07 -0.12 0.06 0.01 -0.07 -0.01 0.04
zero-shot Convo -0.01 -0.03 -0.16 -0.07 -0.23 -0.01 -0.02
News 0.02 -0.07 -0.01 0.06 -0.01 -0.00 -0.03
Science 0.01 -0.01 0.01 0.02 0.01 -0.01 0.01
All 0.01 -0.04 -0.05 0.00 -0.07 -0.01 -0.01
Table 7: Multi-document summarization (MDS) model relative performance when trained and tested on different domains. The results are averaged across all model types. Results are on the test splits of each dataset.
Domain Factuality Accuracy Quality
Train Test Cons. (UE) Rel. (UE) R1 RLSum MoverScore Fluency (UE) Coh. (UE)
ConvoSumm Convo News 0.68 0.41 0.20 0.12 -0.04 0.92 0.45
Convo Science 0.68 0.63 0.23 0.14 0.02 0.91 0.61
News Convo 0.57 0.81 0.21 0.13 -0.02 0.90 0.78
News Science 0.60 0.69 0.24 0.13 0.02 0.93 0.67
Science Convo 0.63 0.38 0.21 0.14 -0.02 0.71 0.40
Science News 0.67 0.39 0.25 0.14 -0.01 0.73 0.43
PRIMERA Convo News 0.78 0.50 0.31 0.16 0.07 0.84 0.65
Convo Science 0.96 0.68 0.25 0.14 0.05 0.83 0.86
News Convo 0.58 0.82 0.18 0.11 -0.01 0.86 0.79
News Science 0.96 0.68 0.25 0.14 0.05 0.83 0.86
Science Convo 0.51 0.60 0.20 0.12 0.01 0.60 0.64
Science News 0.86 0.71 0.35 0.18 0.15 0.65 0.83
SummN Convo News 0.85 0.58 0.24 0.15 0.02 0.93 0.71
Convo Science 0.82 0.62 0.22 0.14 0.01 0.92 0.81
News Convo 0.47 0.82 0.13 0.08 -0.05 0.89 0.80
News Science 0.57 0.62 0.21 0.12 -0.00 0.89 0.60
Science Convo 0.64 0.26 0.10 0.07 -0.16 0.69 0.31
Science News 0.76 0.40 0.33 0.18 0.07 0.83 0.47
Llama 3.1 Convo News 0.83 0.89 0.41 0.22 0.18 0.94 0.93
Convo Science 0.85 0.90 0.26 0.16 0.09 0.93 0.97
News Convo 0.80 0.94 0.21 0.14 0.05 0.93 0.95
News Science 0.86 0.90 0.24 0.15 0.08 0.94 0.96
Science Convo 0.80 0.93 0.21 0.14 0.05 0.93 0.93
Science News 0.82 0.86 0.39 0.23 0.18 0.94 0.91
Llama 3.1 (0-shot) zero-shot Convo 0.79 0.90 0.21 0.15 0.05 0.93 0.92
zero-shot News 0.83 0.84 0.40 0.23 0.19 0.94 0.90
zero-shot Science 0.86 0.90 0.23 0.16 0.08 0.92 0.96
Table 8: Multi-document summarization (MDS) model performance when trained and tested on the different domains. Results are on the test splits of each dataset. Metrics followed by (UE) are from UniEval. For all metrics a greater score is better.
Stage Setting Multi-News ConvoSumm-Reddit Multi-XScience
General
total number of updates 30000 5000 -
warmup updates 500 200 -
learning rate 3e-5 - -
max input sequence 1024 - -
max tokens (in batch) 2048 - -
update frequency 4 - -
patience 2 - -
label smoothing 0.1 - -
dropout 0.1 - -
attention dropout 0.1 - -
weight decay 0.01 - -
optimizer Adam betas=(0.9, 0.999), eps=1e-8 - -
LR scheduler polynomial decay - -
Gradient clipping 0.1 - -
Stage 1
max input sequence 1024 - -
beam 1 - -
length penalty 2 1 -
max length 450 - -
min length 150 0 -
no-repeat-ngram-size 3 0 -
Stage 2
max input sequence 1024 - -
beam 1 - -
length penalty 2 1 -
max length 550 - -
min length 150 0 -
no-repeat-ngram-size 3 0 -
Stage 3
max input sequence 1024 - -
beam 1 - -
length penalty 2 1 -
max length 900 1024 -
min length 400 0 -
no-repeat-ngram-size 3 0 -
Table 9: Training details for SummN on all datasets. Weights initialized from BART-Large-CNN. An entry of “-” indicates the same setting as the previous column.
Setting Multi-News* ConvoSumm-Reddit Multi-XScience*
max input sequence 4096 4096 4096
total number of steps 25K 5K 20K
warmup steps 2.5K 200 2K
optimizer Adam Adam Adam
learning rate 3e-5 3e-5 3e-5
batch size 16 16 16
seed 0 42 0
label smoothing 0.1 0.1 0.1
max length 1024 1024 1024
min length 0 0 0
beam 1 1 1
Table 10: Training details for PRIMERA on all datasets. Models marked with (*) were trained by (Xiao et al., 2022b) and only available training information is filled in. Weights initialized from BART-Large.
Setting Multi-News ConvoSumm-Reddit* Multi-XScience
max input sequence 2048 2048 2048
max tokens (in batch) 2048 - -
update frequency 1 1 1
total number of updates 30000 200 200
warmup updates 500 20 20
learning rate 3e-5 3e-5 3e-5
label smoothing 0.1 0.1 0.1
dropout 0.1 - -
attention dropout 0.1 - -
weight decay 0.01 - -
optimizer Adam betas=(0.9, 0.999), eps=1e-8 - -
LR scheduler polynomial decay - -
Gradient clipping 0.1 - -
max length 1024 1024 1024
min length 0 0 0
beam 1 1 1
length penalty 1 1 1
Table 11: Training details for ConvoSumm on all datasets. Models marked with (*) were trained by (Fabbri et al., 2021a) and only available training information is filled in. Weights initialized from BART-Large. An entry of “-” indicates the same setting as the previous column.
Zero-shot
<|begin_of_text|><|start_header_id|>user<|end_header_id|>
Write a summary that incorporates all of the following documents:
Document: {source document 1}
Document: {source document $n$}
<|eot_id|><|start_header_id|>assistant<|end_header_id|>
In-context Learning
<|begin_of_text|><|start_header_id|>user<|end_header_id|>
Write a summary that incorporates all of the following documents:
Document: {source document 1}
Document: {source document $n$}
<|eot_id|><|start_header_id|>assistant<|end_header_id|>
{target summary}
<|begin_of_text|><|start_header_id|>user<|end_header_id|>
Write a summary that incorporates all of the following documents:
Document: {source document 1}
Document: {source document $n$}
<|eot_id|><|start_header_id|>assistant<|end_header_id|>
Table 12: Prompt templates for Llama 3.1 8B Instruct for the zero-shot (top) and single-shot in-context learning tasks (bottom).

### v2

@@
 We define domain-transfer “failure” as a decrease in factuality, higher deviation from the target, and a general decrease in summary quality.
-In addition to exploring domain transfer for MDS models, we examine potential issues with applying popular summarization metrics out-of-the-box.
+In addition to exploring domain transfer for MDS models, we critically examine the behavior of various automatic evaluation metrics, including their inter-correlations and the performance of LLM-as-a-judge approaches, in these transfer settings.
+Can one size fit all?: Measuring Failure in Multi-Document Summarization Domain Transfer
+Alexandra DeLucia  and Mark Dredze Center for Language and Speech Processing Johns Hopkins University {aadelucia,mdredze}@jhu.edu
 1 Introduction
-Figure 1: Sample generations from SummN trained on Multi-XScience (SummN-Science) and PRIMERA trained on Multi-News (PRIMERA-News) on the ConvoSumm-Reddit dataset (Conversation).
-Relevance is measured automatically with UniEval.
 Multi-document summarization (MDS) is an extended task of summarization, where the goal is to represent the key ideas from multiple documents in a single new document or paragraph.
@@
 While the overarching goal of MDS models is to condense information, models trained on one type of summary (e.g., news) often fail to transfer successfully to other domains (e.g., science) due to the unique nature of how summaries across different domains are written.
-Recent MDS system surveys compare different models either on a single domain (e.g., conversation/dialogue (Zhang et al., 2021; Mullick et al., 2024)) or focus on MDS corpora biases (Wolhandler et al., 2022; Dey et al., 2020).
+Domain-Transfer Example
+Ground-Truth A couple of the responses to this post are either satirical or sarcastic. ...
+News-based Might you be a little nervous about playing a video game for free? you’re not alone. ...
+Science-based @cite @cite have both mentioned that the reason for the lack of a sub is due to Microsoft’s refusal to allow sub-free games on Xbox. ...
+Figure 1: Excerpt generations two types of multi-document summarization models (SummN and ConvoSumm), trained on different datasets (Multi-News+ and Multi-XScience), and both evaluated on a conversation summarization task (ConvoSumm-Reddit).
+The ROUGE-LSum of the generations to the Ground-Truth are 0.13 and 0.12, respectively, and UniEval-Relevancy scores of 0.78 and 0.3, respectively.
+Recent MDS system surveys compare different models either on a single domain (e.g., conversation/dialogue (Zhang et al., 2021; Mullick et al., 2024) or focus on MDS corpora biases (Wolhandler et al., 2022a; Dey et al., 2020).
 While Zhang et al.
@@
 (2021a) a zero-shot model from news to conversation had lower performance than on conversation alone.
-An example is in Figure 1, where two models, one trained on Multi-News (Fabbri et al., 2019) and the other on Multi-XScience (Lu et al., 2020) are evaluated on a sample from ConvoSumm-Reddit (Fabbri et al., 2021a).
-The difference in writing style is immediately apparent: the news-trained model draws out quotes, the science-trained model inserts citations, and both are completely different from the expected style of the Reddit conversation summary.
+An example is in Figure˜1, where two models, one trained on Multi-News+ (Choi et al., 2024; Fabbri et al., 2019) and the other on Multi-XScience (Lu et al., 2020) are evaluated on a sample from ConvoSumm-Reddit (Fabbri et al., 2021a).
+The difference in writing style is immediately apparent: the news-trained model has a blogger-like style, the science-trained model inserts citations, and both are completely different from the expected style of the Reddit conversation summary.
 The difference in style is also captured by similarity metrics ROUGE (Lin, 2004) and UniEval-Relevance (Zhong et al., 2022)–while the ROUGE scores, which capture word overlap, only differ by a few points, the semantic-based, trained Relevance scores vary widely.
@@
 Similarly, summaries for other domains also vary in incorporated information and style.
-Figure 1 highlights the three axes of domain transfer for MDS systems: 1) domain of the training data, 2) type of model, and 3) evaluation metrics.
+Figure˜1 highlights the three axes of domain transfer for MDS systems: 1) domain of the training data, 2) type of model, and 3) evaluation metrics.
 While there are popular benchmarks for summarization, including the notable SummEval (Fabbri et al., 2021b), few are for multi-document summarization, and most datasets focus on the news domain.
@@
 We find that: [^1] All code and models will be released upon publication.
-  • Smaller language models (BART-size) hold more stylistic bias with their training data;
-  • Within News, Science, and Conversation, there is no clear “best” domain for general transfer;
-  • Lexical and semantic similarity metrics both fail to capture summary quality;
-  • Current non-reference-based metrics do not scale well to the multi-document setting;
-  • Trained evaluators have bias and penalize out-of-domain datasets;
+  • LLMs (e.g., Llama 3.1) demonstrate robust quality and factuality in domain transfer with minimal performance change, and their outputs are less influenced by few-shot examples compared to the stronger stylistic bias observed in smaller, fine-tuned models;
+  • While no single source domain (News, Science, or Conversation) proves universally superior for general transfer, the stylistic nature of summaries learned during training significantly impacts cross-domain performance;
+  • Lexical similarity metrics (e.g., ROUGE, MoverScore) show weak correlation with semantic relevance (e.g., UniEval-Relevance) and other quality dimensions;
+  • Source-dependent non-reference metrics face scaling challenges moving from single-document summarization to multi-;
+  • Trained multi-dimensional evaluators exhibit intra- and inter-correlations with each other and within dimensions;
+  • The number of source documents surprisingly shows weak and inconsistent correlations with factuality and coherence metrics in domain-transfer settings across different model architectures;
 Figure 2: Summary of all the domains (datasets), multi-document summarization (MDS) model types, and evaluation metrics analyzed in this work.
+The “Relevance” metric is in both Ground-Truth Similarity and Overall Quality groups because the UniEval (UE) and Prometheus (P) definitions are different (see Figures˜10 and 11 for details).
 2 Domain Transfer of Multi-Document Summarization
@@
 The main dimensions of our analyses are domain (i.e., dataset), model type, and evaluation metrics.
-See Figure 2 for summary.
+See Figure˜2 for summary.
 2.1 Summarization Dataset Domains
@@
 News.
-Multi-News (Fabbri et al., 2019) is a news-based MDS dataset featuring 56,216 human-written summaries from Newser.222newser.com
+Multi-News+ (Choi et al., 2024), an automatically cleaned version of Multi-News (Fabbri et al., 2019), is a news-based MDS dataset featuring human-written summaries from Newser.222newser.com.
+Multi-News+ contains 44,668/5,585/5,584 entries in the train/validation/test sets.
 Science.
@@
 Since the paper dump was from pre-2020, we assume all writing on arXiv is human-generated.
-2.2 Multi-Document Summarization Model Types
-Current high-performing fine-tuned abstractive summarization methods can be grouped into four approaches: summary-specific pre-training (Zhang et al., 2020a; Xiao et al., 2022a; Peper et al., 2024), incremental summarization (“chunk-then-summarize”) (Zhang et al., 2022), extracting claims or viewpoints (“extract-then-summarize”) (Fabbri et al., 2021a; Ernst et al., 2022; Ouyang et al., 2023), and fine-tuning or few-shot summarization with large pre-trained models (Mullick et al., 2024).
+2.2 Multi-Document Summarization Models
+Current high-performing fine-tuned abstractive summarization methods can be grouped into four approaches: summary-specific pre-training (Zhang et al., 2020a; Xiao et al., 2022; Peper et al., 2024), incremental summarization (“chunk-then-summarize”) (Zhang et al., 2022), extracting claims or viewpoints (“extract-then-summarize”) (Fabbri et al., 2021a; Ernst et al., 2022; Ouyang et al., 2023), and fine-tuning or few-shot summarization with large pre-trained models (Mullick et al., 2024).
 These groupings have been used in prior work (Mullick et al., 2024).
 Modified Pretraining.
-Prior work has shown that fine-tuning a language model specifically pre-trained with summary-focused objectives improves downstream performance over non-specialized models (Xiao et al., 2022b; Zhang et al., 2020a; Peper et al., 2024).
+Prior work has shown that fine-tuning a language model specifically pre-trained with summary-focused objectives improves downstream performance over non-specialized models (Xiao et al., 2022; Zhang et al., 2020a; Peper et al., 2024).
 Zhang et al.
-(2020a) first introduced this concept of summary-specific pre-training, and it was later augmented to the multi-document case (Xiao et al., 2022b; Peper et al., 2024).
+(2020a) first introduced this concept of summary-specific pre-training, and it was later augmented to the multi-document case (Xiao et al., 2022; Peper et al., 2024).
 We select PRIMERA to represent this group.
-PRIMERA is Longformer-Encoder-Decoder model initialized from BART-Large (Beltagy et al., 2020) and pre-trained with an Entity Pyramid objective on a large news corpora (Xiao et al., 2022b).
+PRIMERA is Longformer-Encoder-Decoder model initialized from BART-Large (Beltagy et al., 2020) and pre-trained with an Entity Pyramid objective on a large news corpora (Xiao et al., 2022).
 While PELMS (Peper et al., 2024) adjusts the pre-training strategy to improve upon PRIMERA, we omit comparison to this model because it also uses a data filtering step, which overlaps with the extract-then-summarize approach.
@@
 The complement of ROUGE is BLEU (Papineni et al., 2002), which measures precision instead of recall of generated summaries.
-Other metrics in the spirit of ROUGE/BLEU are semantic-based instead of lexical, which do not penalize generations with similar meanings but different word choices.
+Other metrics in the spirit of ROUGE/BLEU are semantic-based instead of lexical, which are more robust to passages with similar meanings but different word choices.
 The most popular version is BERTScore (Zhang et al., 2020b), which maps tokens from the generation to the reference before calculating overlap.
-An alternative is MoverScore, which “optimized” version of BERTScore (Zhang et al., 2020b) which uses soft token alignments from the hypothesis to the reference instead of a hard 1:1 mapping, allowing for a greater reward for semantically similar generated text.
-Other similarity metrics are sub-“dimensions” of a trained multi-dimensional evaluator, like UniEval.
+An alternative is MoverScore, an “optimized” version of BERTScore which uses soft token alignments from the hypothesis to the reference instead of a hard 1:1 mapping, allowing for a greater reward for semantically similar generated text.
+Other similarity metrics are sub-“dimensions” of a trained multi-dimensional evaluator, like UniEval (Zhong et al., 2022).
 UniEval is a T5-Large model fine-tuned on synthetically augmented datasets for positive and negative examples of different “dimensions” of summary attributes (fluency, relevance, coherence, consistency) (Zhong et al., 2022).
@@
 Overall Quality.
-Not all desired qualities of a summary depend on a reference, some are intrinsic or rely on the source documents (e.g., fluency).
+Not all desired summary qualities depend on a reference; some are intrinsic or rely on the source documents.
+Intrinsic metrics include fluency (i.e., well-written and grammatical), coherence (i.e., overall flow), and the summarization-specific metrics of compression ratio (i.e., length of source documents compared to length of the summary) and source coverage (i.e., how well the summary represents the source documents).
 UniEval measures both fluency and coherence.
 UniEval-Fluency is based solely on the generation, and UniEval-Coherence measures the coherence of the generation with respect to the source.
+Similar to UniEval, Prometheus (Kim et al., 2024) is another multi-dimensional evaluator, but is a general-purpose feedback model that can judge any model output according to a provided rubric.
+In this respect, Prometheus is essentially a distilled “LLM-as-a-Judge” model (Zheng et al., 2023).
+For a more up-to-date automatic evaluation method, we use Prometheus to grade the summaries on the following criteria, drawn from SummEval (Fabbri et al., 2021b): Relevance, Coherence, and Fluency.
+An important note is that despite the same metric names, the definitions from SummEval differ slightly from those in UniEval.
+To measure compression ratio, we include both Compression-Sentence, which compares the ratio of the number of sentences in the summary to those in the source documents, and Compression-Word, which compares the ratio of the number of words in the summary to those in the source documents (Koh et al., 2022).
+For source coverage (Wolhandler et al., 2022a), we developed a prompt for Prometheus, Prometheus-Coverage.
 Factuality.
@@
 While objective factuality is important for the safety of LLMs (e.g., knowing that boiling water is hot), for MDS we typically assume that the source documents are factually correct and that a summary is only factual if its claims match the claims from the source.
-We measure summary factuality with UniEval-Consistency, a single-dimensional evaluator trained to measure the factual consistency of generated text to the source (as opposed to continuously trained as with UniEval-Fluency and UniEval-Coherence).
-A popular framework for factuality analysis is decompose-then-verify (Min et al., 2023; Jiang et al., 2024), and we leave this analysis to future work.
-We include ROUGE, MoverScore, and multiple dimensions of UniEval in our analysis as representations for similarity to ground truth, factuality, and overall quality metrics.
-While BARTScore (Yuan et al., 2021) measures many of the same dimensions as UniEval, we chose UniEval because it is better correlated with human evaluation (Zhong et al., 2022) and the potential bias of BARTScore to model generations, since most of the models are BART-based (“self bias”) (He et al., 2023).
-Another evaluation for summaries is source coverage, or how much of the source documents are represented in the summaries (Wolhandler et al., 2022).
-We leave this evaluation to future work.
+This definition of “factually” is better described as grounding in the source text.
+We measure summary factuality with UniEval-Consistency, a single-dimensional evaluator trained to measure the factual consistency of generated text to the source (as opposed to continuously trained as with UniEval-Fluency and UniEval-Coherence), Prometheus-Consistency, and Prometheus-Factuality.
+The Prometheus-Consistency prompt is drawn from SummEval, and the Prometheus-Factuality prompt is from the Prometheus repository.
+[…]
 3 Experimental Setup
 Our goal is to explore the intricacies of domain transfer with multi-document summarization models and their evaluation methods.
-The main dimensions of our analyses are different datasets (i.e., domains), models, and evaluation metrics (Figure 2).
-We chose 3 dataset domains (News, Science, Conversation), 4 model types (end-to-end/direct), chunk-then-summarize, extract-then-summarize, and pre-trained LLMs), and 7 metrics for a total of 84 points of analysis (Section 2).
+The main dimensions of our analyses are different datasets (i.e., domains), models, and evaluation metrics (Figure˜2).
+We chose 3 dataset domains (News, Science, Conversation), 4 model types (end-to-end/direct), chunk-then-summarize, extract-then-summarize, and pre-trained LLMs), and 15 metrics for a total of 180 points of analysis (Section˜2).
 3.1 Model Training
-The hyperparameters for each model and dataset combination are in Appendix B.
-Each model was trained on each dataset, and we refer to them as ModelName-Domain, e.g., ConvoSumm fine-tuned on Multi-News is ConvoSumm-News.
+The hyperparameters for each model and dataset combination are in Appendix˜B.
+Each model was trained on each dataset, and we refer to them as ModelName-Domain, e.g., ConvoSumm fine-tuned on Multi-News+ is ConvoSumm-News.
 PRIMERA.
@@
 To handle examples with source documents that are too long for the context window, we follow the author’s code and evenly truncate the documents to fit.
-We used the authors’ released models of PRIMERA fine-tuned on Multi-XScience and MultiNews and trained PRIMERA-ConvoSumm-Reddit with the released author code.
+We used the authors’ released PRIMERA-MultiXScience model and trained PRIMERA-Convo and PRIMERA-News with the released code.
+[^3] While there is a PRIMERA model trained on Multi-News, we trained a new one on Multi-News+.
 SummN.
 We use a 3 course-stage process for all datasets.
-A separate model is trained for each coarse stage, leading to $N+1$, or $4$ fine-tuned BART models for each training dataset.
+A separate model is trained for each coarse stage, leading to $N+1$, or 444 fine-tuned BART models for each training dataset.
 An important note is due to the setup of summarizing chunks, there are no too-long context issues as with direct models like PRIMERA and all source documents are considered by the model.
@@
 Unlike the other models, we do not fine-tune Llama 3.1 8B-Instruct.
-[^3] https://huggingface.co/meta-llama/Llama-3.1-8B-Instruct To compare to the domain-transfer experiments, we provide an example from the training set to the model, i.e., for News-Convo, an example from Multi-News is provided to Llama and then the model is prompted to complete an example from ConvoSumm.
-We also include a zero-shot version which does not provide a complete example to Llama.
+[^4] https://huggingface.co/meta-llama/Llama-3.1-8B-Instruct To compare to the domain-transfer experiments, we provide an example from the training set to the model, i.e., for News-Convo, two randomly selected examples from Multi-News+ is provided to Llama and then the model is prompted to summarize an example from ConvoSumm.
+We also include a zero-shot version which does not provide a complete example to Llama, only the instruction.
 3.2 Evaluation Details
@@
 For the other evaluation metrics, MoverScore and multiple dimensions from UniEval, we use the trained models as they are without modification.
-Factuality Ground-Truth Similarity Quality
-Cons. (UE) Rel. (UE) R1 RLSum MoverScore Fluency (UE) Coh. (UE)
-Conversation 0.68 0.75 0.30 0.20 0.07 0.91 0.72
-Avg. $\Delta$ +18% -13% -9% -19% -31% -1% +4%
-Ref. 0.71 0.99 1.00 1.00 1.00 0.90 0.73
-News 0.72 0.78 0.43 0.22 0.20 0.91 0.79
-Avg. $\Delta$7% +2% -51% -41% -92% -2% +1%
-Ref. 0.72 0.99 1.00 1.00 1.00 0.92 0.80
-Science 0.78 0.63 0.24 0.15 0.03 0.78 0.64
-Avg. $\Delta$7% -12% +6% +1% -7% -1% +4%
-Ref. 0.55 0.99 1.00 1.00 1.00 0.66 0.43
-Table 1: Multi-document summarization (MDS) model performance and relative performance aggregated by domain when trained and tested on the same domain (top) and different domains ($\Delta$), respectively.
-Results are aggregated across all the test splits of each dataset.
-A positive $\Delta$ shows a performance gain and a negative $\Delta$ is a performance loss as compared to the model’s scores on the in-domain data.
-A relative change of over $1$ is possible with MoverScore because of its $[-1,1]$ range.
-Detailed results are in Table 7.
-“Ref” refers to the results when running the metrics on the ground-truth data.
+For evaluation with Prometheus, we use Prometheus2-BGB-8x7B  [^6] https://huggingface.co/prometheus-eval/prometheus-bgb-8x7b-v2.0 as an absolute judge with dimension-specific rubrics we based off the definitions from SummEval (Fabbri et al., 2021b).
+See Appendix˜C for details.
+Due to the intense compute required for Prometheus, we sampled 100 items from each dataset for all Prometheus-based metrics.
+ConvoSumm PRIMERA SummN Llama 3.1
+Llama 3.1
+(0-shot)
+Avg. $\Delta$ Avg. $\Delta$ Avg. $\Delta$ Avg. $\Delta$ Avg. $\Delta$
+Factuality Consistency (UE) 0.67 -2% 0.82 0% 0.65 +3% 0.86 0% 0%
+Factuality (P) 2.04 -16% 1.79 +3% 2.02 -6% 3.23 0% 0%
+Consistency (P) 2.14 -15% 1.89 +1% 2.08 -7% 3.40 +1% 0%
+Ground-Truth Similarity Relevancy (UE) 0.66 -3% 0.48 +6% 0.66 -19% 0.92 0% 0%
+ROUGE-1 0.35 -27% 0.27 +12% 0.35 -30% 0.29 +10% 0%
+ROUGE-LSum 0.20 -28% 0.14 +6% 0.20 -28% 0.17 +4% +1%
+MoverScore 0.10 -86% 0.06 -332% 0.10 -88% 0.10 +35% -2%
+Quality Relevancy (P) 2.36 -15% 2.18 +3% 2.33 -6% 3.53 +1% +1%
+Fluency (UE) 0.79 -5% 0.69 +2% 0.76 0% 0.94 0% 0%
+Fluency (P) 2.30 -21% 1.67 +9% 2.26 -5% 3.97 0% 0%
+Coherence (UE) 0.60 -14% 0.70 +2% 0.63 -2% 0.95 0% 0%
+Coherence (P) 2.31 -20% 1.63 +8% 2.31 -11% 3.95 0% 0%
+Comp.-Sent 0.71 +7% 0.66 +1% 0.69 +9% 0.64 +4% 0%
+Comp.-Word 0.74 +1% 0.55 +2% 0.70 -1% 0.55 +2% 0%
+Coverage (P) 1.93 -12% 1.83 +3% 1.95 -6% 3.03 +1% 0%
+Table 1: Multi-document summarization (MDS) model performance and relative performance aggregated by model when trained and tested on the same domain (left) and different domains (right, $\Delta$), respectively.
+A relative change of over 111 is possible with MoverScore because of its [1,1][-1,1][ - 1 , 1 ] range.
+Llama (0-shot) is compared to Llama with in-domain examples.
 4 Results
 We analyze the results in the framework of our three axes: dataset domain, model type, and evaluation metric.
-For ease of analysis, we show the relative performance of each model when evaluated on a new domain in Tables 2 and 1.
-The relative performance is the percentage gain or loss of a model’s performance when compared to the performance on the in-domain data.
-The zero-shot baseline with Llama 3.1 is compared to the in-context Llama 3.1 performance for each dataset (i.e., no examples instead of one example).
-See Appendix C for the detailed per-model and dataset results.
-4.1 Domain Generalizability
-When aggregated by dataset over the different model types, clear patterns emerge from Table 1.
-Science summaries have the lowest general quality.
-The quality is lowest in the Science domain compared to Conversation and News according to Fluency and Coherence metrics, even before evaluating Science-trained MDS models on other domains, $0.78$ and $0.64$, respectively.
-These scores are $10+$ points lower than those of News ($0.91$/$0.79$) and Conversation ($0.91$/$0.72$).
-Fluency is affected the least by domain transfer.
-The fluency scores are the most stable, i.e., have the least change in performance (Avg $\Delta$) compared to the other metrics.
-While on average there is a slight decline in fluency, as indicated by the negative change, the decline is negligible in the $(-0.02,-0.01)\%$ range.
-Fluency does not appear to be an issue with these models, mostly likely due their size (BART-Large and Llama 8B).
-Factuality Ground-Truth Similarity Quality
-Cons. (UE) Rel. (UE) R1 RLSum MoverScore Fluency (UE) Coh. (UE)
-ConvoSumm 0.65 0.62 0.32 0.18 0.08 0.83 0.57
-Avg. $\Delta$2% -11% -28% -26% -152% +3% -4%
-PRIMERA 0.78 0.74 0.35 0.20 0.12 0.89 0.78
-Avg. $\Delta$ +3% -10% -19% -24% -16% -14% -1%
-SummN 0.66 0.60 0.33 0.19 0.08 0.81 0.56
-Avg. $\Delta$ +4% -9% -32% -33% -31% +7% +18%
-Llama 3.1 0.82 0.91 0.29 0.18 0.11 0.94 0.94
-Avg. $\Delta$ +1% -1% +7% +4% +25% -0.00% +0.00
-Llama 3.1 (0-shot) +1% -4% -5% +0% -7% -1% -1%
-Table 2: Multi-document summarization (MDS) model performance and relative performance aggregated by model when trained and tested on the same domain (top) and different domains ($\Delta$), respectively.
-Results are aggregated across all the test splits of each dataset.
+[^7] The domain-transfer analysis is in Appendix D.
+See Appendix˜E for the detailed per-model and dataset results.
+To study evaluation metrics, we analyze the performance of the ground-truth dataset.
+Complete summary examples from each model are in Appendix Figures˜14, 16, 13, 12 and 15.
+4.1 MDS Model Generalizability
+To study trends specific to each model type, we aggregated the metrics across all the datasets (Table˜1).
+We also look at correlations between metrics in Appendix Figure˜8.
+Instead of judging performance by the raw metrics, we look at the relative change in performance from the in-domain setting to the out-of-domain setting (e.g., training on ConvoSumm-Reddit and testing on Multi-XScience).
+The zero-shot baseline with Llama 3.1 is compared to the in-context Llama 3.1 performance for each dataset.
 A positive $\Delta$ shows a performance gain and a negative $\Delta$ is a performance loss as compared to the model’s scores on the in-domain data.
-A relative change of over $1$ is possible with MoverScore because of its $[-1,1]$ range.
-LLama zero-shot is compared to Llama with in-domain examples.
-Detailed results are in Tables 5 and 6.
-There is no clear general-purpose domain for transferring.
-Each domain has different benefits and cons to train on for transferring, as indicated by the varied relative performance scores ($\Delta$).
-Transferring from Conversation to News and Science can lead to greatly increased factuality, as measured by UniEval-Consistency ($+18\%$), but also a decline in Ground-Truth Similarity ($-31\%$ to $-9\%$).
-Transferring from News appears to be the worst according to the similarity metrics, over $-92\%$ according to MoverScore.
-The Science-trained models appear to be the most neutral overall, ranging from $-12\%$ to $+6\%$ across all metrics, but claiming that it is the best domain for general MDS model training is difficult because the model scores started lower than the other domains to begin with.
-News-trained models have the worst similarity when transferring to non-News data.
-There is a steep decline in performance when News-trained models are evaluated on Conversation and Science as indicated by ROUGE ($-51\%$ R1, $-41\%$ RL-Sum) and MoverScore ($-92\%$).
-However, the Relevance score is high ($+2\%$).
-The discrepancy in word-mapping metrics (lexical with ROUGE and semantic with MoverScore) and overall similarity (Relevance) show that a summary can still incorporate information from the ground-truth without having an overlap in words.
-Conversation and News data have the most unique styles.
-The ConvoSumm-Reddit example from Figure 1 shows the stark difference in style between the types of summaries in ConvoSumm-Reddit and those from Multi-News.
-This difference in style explains the performance loss from Conversation and News domains, as the summary wording would be completely different and penalized by the ground-truth similarity metrics ($-31$ to $-9\%$ and $-92$ to $+2\%$, respectively).
-4.2 MDS Model Transfer
-Grouping the domain transfer results by model type across all dataset domains reveals even more performance trends (Table 2).
-Pre-trained LLM outperforms all fine-tuned smaller models.
-For all metrics besides similarity, Llama 3.1 outperforms the other models concerning factuality and quality.
-This trend holds during domain transfer, with a performance difference ranging from $-1$ to $+25\%$ across all metrics.
-The second-best performing model across all metrics, including accuracy, is PRIMERA, indicating an advantage of directly fine-tuning a long-context model (4K context window).
+Llama 3.1 outperforms all fine-tuned smaller models in quality and factuality metrics.
+Across all datasets in the in-domain setting, Llama generates the most factual and highest-quality summaries according to the automatic metrics.
+This trend holds during domain transfer, with minimal change in performance (+1+1+ 1 to +4+4+ 4).
+The exception to the higher performance trend is with the compression metrics and lexical-similarity metrics.
+Llama has the lowest compression, similar to PRIMERA, as summaries are only $\%5$ compressed.
 Zero-shot performance is the same as in-context learning with Llama 3.1.
 Surprisingly, Llama has a similar performance at generating summaries with (in-context learning, ICL) or without being shown an example (zero-shot).
+The Llama-generated summaries change very little when shown examples versus when no examples are provided (Appendix Figures˜15 and 16).
 This indicates a strong bias to its pre-training versus “learning” from the example.
 These results could differ with the addition of more examples (at the expense of a smaller context window available for the source documents) or a larger model.
-ConvoSumm claim extraction is ineffective on Science.
-The first step in the extract-then-summarize model ConvoSumm is to filter out sentences that are not identified as claims by a small, fine-tuned model.
-This step removes $35\%$ of sentences on average from each ConvoSumm-Reddit example, $51\%$ from Multi-News, but less than $1\%$ from Multi-XScience.
-This is intuitive since there would be more filler sentences in conversations and news articles, but less so in scientific publications.
+Not all domains benefit from Extract-then-Summarize models.
+[…]
 Chunk-then-summarize is the only approach guaranteed to fit all source documents.
 With all models, including Llama 3.1 with a 131K context window, there were issues fitting all source documents for every example comfortably in the context windows.
-As mentioned in Section 3, PRIMERA “handles” this issue by evenly dividing the context window among the source documents and truncating the ends.
+As mentioned in Section˜3, PRIMERA “handles” this issue by evenly dividing the context window among the source documents and truncating the ends.
 The assumption that the beginning of documents is more important than the end does not hold for all domains and has only been identified in News (Dey et al., 2020).
@@
 Llama only had issues with 3 samples from Multi-News.
+The number of source documents does not have strong correlations with other metrics.
+Surprisingly, the number of source documents is weakly and inconsistently correlated with factuality and coherence metrics (0.1-0.1- 0.1 to 0.5-0.5- 0.5).
+This is different than expected, as model performance is expected to decrease with longer context and more information to process.
+All models are fluent.
+Unsurprisingly, all models have high fluency scores (0.70+0.70+0.70 +) since all models are of size BART-Large and above.
+While fluency is still an important attribute to check, it might not be as important as it once was.
+Word-based similarity metrics are not correlated with Relevancy.
+While UniEval-Relevancy measures the similarity between a generated and reference summary (“Is this summary relevant to the reference?”), it does not correlate with the semantic and lexical metrics like MoverScore and ROUGE (0.1-0.1- 0.1 to 0.10.10.1).
+Whereas MoverScore and ROUGE are highly correlated (0.60.60.6 to 0.80.80.8).
 4.2 Evaluation of Evaluation
-Word-based similarity scores are low across all models.
-For all models, the generated summaries are not very similar to the ground-truth reference according to the lexical and semantic similarity metrics ROUGE and MoverScore.
-The model with the highest accuracy ($0.20$ ROUGE_L and $0.12$ MoverScore) was PRIMERA, higher than the much larger LM, Llama-8B.
-For lexical-based metrics like ROUGE, the low similarity can be explained through the model’s use of synonyms, which would not be captured through token-based recall.
-However, the low MoverScores can be explained by (1) the penalization of generations longer or more detailed than the reference and (2) the lack of a normalizing constant as provided by BERTScore (Zhang et al., 2020b).
-[^5] Zhang et al.
-(2020b) use a learned constant to normalize the scores to $[0,1]$ for easier readability.
-Still, low scores across both accuracy metrics indicate that the generated summaries differ from their human-written counterparts.
-Despite overall low accuracy scores, all models have the highest ROUGE/MoverScore performance on Multi-News (News), indicating the summaries might not be as abstract as the other datasets.
-Surprisingly, the Relevance score which compares the reference and generation as a whole instead of token mappings, shows higher similarity, especially for Llama ($0.91$ average).
-Accuracy is affected the most.
-The reference-based metric, accuracy, shows the most extreme gains and losses compared to the other metrics.
-This is intuitive because a model’s summary can still be good but not be written in the same manner, or style, as the reference, which can happen if a model is biased towards the writing style of its training data.
-The zero-shot baseline, which is only prompted and not given any styles, shows the least variance in accuracy compared to the other models.
-All models are fluent.
-Unsurprisingly, all models have high fluency scores ($0.80+$) since all models are of size BART-Large and above.
-While fluency is still an important attribute to check, it might not be as important as it once was.
+To isolate patterns of the evaluation metrics, we look at the automated metric values on the ground-truth reference datasets (Appendix Table˜3) and their correlations (Figure˜3).
+See Appendix˜E for detailed correlation figures.
+Figure 3: Pearson correlations across the ground-truth for all evaluated datasets.
+The black squares section off intra-correlations within Prometheus and UniEval, and the red square highlights the correlations between them.
+UniEval and Prometheus have a weak to moderate correlation.
+Over the three datasets, the UniEval and Prometheus metrics have weak to moderate correlations (0.30.30.3 to 0.50.50.5).
+While Prometheus is $60x$ larger than UniEval (T5-Large), a stronger positive correlation is expected when they are evaluating the same quality (e.g., fluency).
+The extreme outlier is UniEval-Relevance, which has poor to no correlation with any other metric (0.00.00.0 to 0.20.20.2).
+Metric names are not consistent across evaluators and benchmarks.
+The definition of “relevance” differs between UniEval and SummEval, “Is this summary relevant to the reference?” versus “How well does the summary capture the key points of the documents?”.
+This discrepancy explains the 0.00.00.0 correlation between UniEval-Relevance and Prometheus-Relevance.
+The other metrics have similar definitions and are more correlated, as noted above.
+See Appendix Figures˜10 and 11 for dimension definitions.
+Multi-dimensional evaluators are slightly intra-correlated.
+Within Prometheus, the dimensions have slight (0.40.40.4) to moderate (0.70.70.7) correlations with each other, specifically coherence, consistency, and factuality (0.60.60.6 to 0.70.70.7).
+The UniEval dimensions are less intra-correlated, (0.00.00.0 to 0.40.40.4).
+This discrepancy in intra-correlation could indicate that Prometheus has difficulty evaluating along a single dimension and instead allows various quality aspects to impact the assigned score.
+This issue persisted despite adding instructions only to evaluate the desired dimension (see Appendix˜C).
 Non-reference metrics do not scale well to the multi-document setting.
-Evaluating for attributes that depend on the source documents (Consistency and Coherence) is most affected by the long-document nature of multi-document summarization.
+Evaluating for attributes that depend on the source documents (e.g., Consistency) is most affected by the long-document nature of multi-document summarization.
 The options are either to 1) truncate the documents, which can greatly alter the scores depending on which documents are included, or 2) calculate the metric for every source document and aggregate the scores.
-More work is needed to determine whether truncating the source documents for UniEval over- or under-estimates model performance.
-Another option is to turn to large-context models like GPT-4 to be a “judge” for specific criteria (Kim et al., 2024; Vu et al., 2024).
-This use of LLMs, which typically can fit 10K+ tokens in their context windows, is a potential solution to this scaling issue.
-The reference-based metrics, i.e., similarity metrics that only compare the ground-truth and generated summaries, are not affected by this issue.
-Trained metrics can penalize out-of-training data.
-Just as summarization models can have lower performance on out-of-domain data, so can trained evaluation models like UniEval.
-As noted before, the fluency and coherence scores on Science are $0.10$ points lower than on News and Conversation.
-This could be an issue with the summarization model training, but the fluency and coherence scores are also lower on the ground-truth summaries, $0.30+$ points lower (“Ref.” in Table 1), which should not be the case as all summaries are human-written and should have high scores.
-An explanation is that UniEval was trained on a synthetically augmented version of news data, CNN/DailyMail, which is very different from Scientific text.
+More work is needed to determine whether truncating or aggregating per-document scores over- or under-estimates model performance.
+As for the now-popular “LLM-as-a-Judge” models, like Prometheus, while they have less of a context-window issue, they have a higher computational cost.
+The reference-based metrics (primarily similarity-based) are not affected by scale because they only compare the ground-truth and generated summaries.
 5 Discussion and Conclusion
 In most work that surveys or proposes multi-document summarization (MDS) systems, there is little overlap between the analysis of model performance and the shortcomings of popularly used summarization evaluation metrics.
-We bridged this gap through experiments and analysis of not only the zero-shot domain transfer ability of MDS systems but also their evaluation.
-We analyzed three components of a multi-document summarization (MDS) pipeline for domain transfer: 1) the domain of the training data, 2) the model type, and 3) the common evaluation metrics.
-We found important facets to consider in each component.
-For domain transfer, the training and test domains are important for model performance, regardless of the model type.
-While we did not find a clear trend between Conversation, News, and Science, it was apparent that the transfer was highly affected by the style of the summaries, albeit more than the domain itself.
-Summaries from different domains but written in the same style (e.g., high-level vs low-level) could transfer better with higher similarity to the reference.
-For different approaches to MDS, a model that didn’t hold a strong bias towards the style of its training (or example) data is ideal, as is the case with a large pre-trained decoder model, like Llama 3.1.
-If the domain styles differ greatly, a zero- or few-shot approach with less bias is ideal.
-Notable issues we uncovered with evaluation metrics that prevented the fair evaluation of summaries were domain-specific penalizations, mostly for Science, questions with scaling to the MDS case, and the inability of ground-truth similarity metrics to adjust for style change (ROUGE and MoverScore).
-When evaluating an MDS model on a new domain, word-similarity metrics judge style more than content similarity and trained evaluators could over- or under-judge model performance due to limited information in the context window.
+We bridged this gap through experiments and analysis of not only the zero-shot domain transfer ability of MDS systems but also a critical examination of their evaluation methodologies.
+We analyzed three components of the MDS pipeline for domain transfer: 1) the domain of the training data, 2) the model type, and 3) the common evaluation metrics, including their inter-correlations and the nuances of LLM-as-a-judge approaches.
+In conclusion, this work underscores the intricate relationship between MDS model architectures, the stylistic properties of data domains, and the evaluation metrics used to gauge performance.
+For models, a reduced stylistic bias and the ability to handle extensive context are beneficial for domain transfer.
+For evaluation, our findings strongly advocate for a multi-faceted approach, exercising caution with lexical metrics for advanced models, and a keen awareness of the specific definitions, potential biases, scaling limitations, and inter-correlation patterns of even sophisticated trained evaluators.
+Future research should continue to develop more robust and reliable evaluation strategies that are sensitive to the nuances of abstractive summarization across diverse domains and model capabilities.
 Limitations
@@
 We leave these tasks to future work or the incorporation of an additional step for model safety.
+AI Assistants were used in the preparation of this work, specifically for paper revision, table formatting, and figure editing.
+Acknowledgements
 Appendix A Dataset Details
-We include three MDS datasets: Multi-News (Fabbri et al., 2019), Multi-XScience (Lu et al., 2020)), and ConvoSumm-Reddit (Fabbri et al., 2021a).
+We include three MDS datasets: Multi-News+ (Choi et al., 2024), Multi-XScience (Lu et al., 2020)), and ConvoSumm-Reddit (Fabbri et al., 2021a).
 We keep the original train, validation, and test splits as the original papers for all datasets.
-See Table 3 for an example from each dataset and Table 4 for summary statistics.
-Multi-News.
-For Multi-News we used the “cleaned” version provided by the authors [^6] https://github.com/Alex-Fabbri/Multi-News/issues/36 with 44964/5621/5621 examples for train, validation, and test sets, respectively.
-We keep the NEWLINE_CHAR symbol in the documents.
+See Figures˜6, 4 and 5 for an example from each dataset and Table˜2 for summary statistics.
+Multi-News+.
+This dataset has 37,057/4,576/4,648 examples for train, validation, and test sets, respectively. We keep the NEWLINE_CHAR symbol in the documents.
 Multi-XScience.
@@
 Unless specified otherwise, each model was trained on a single Nvidia A100 GPU with 80GB of memory.
-The training hyperparameters for PRIMERA, SummN, and ConvoSumm-arg-filtered are in Tables 10, 9 and 11, respectively.
+The training hyperparameters for PRIMERA, SummN, and ConvoSumm-arg-filtered are in Tables˜7, 6 and 8, respectively.
 Below are the prompts and details for running Llama for the zero-shot and in-context 1-shot comparisons.
@@
 Due to the large size of Llama 8B, we hosted it as a service with vLLM on 4 Nvidia A100 80GB GPUs (Kwon et al., 2023).
-[^7] https://docs.vllm.ai We followed model examples on how to interact for the summarization task (Table 12).
-Appendix C Raw Results
-The domain-transfer results in Section 4.2 show the summarized relative performance of the models.
-The detailed performance across domains and models are in Tables 7 and 6, respectively.
-The raw, un-normalized numbers are shown in Table 8.
-Dataset Source Documents Summary
-Multi-News Whether a sign of a good read; or a comment on the ’pulp’ nature of some genres of fiction, the Oxfam second-hand book charts have remained in The Da Vinci Code author’s favour for the past four years.
-NEWLINE_CHAR NEWLINE_CHAR Dan Brown has topped Oxfam’s ’most donated’ list again, his fourth consecutive year …
-A charity shop is urging people to stop donating The Da Vinci Code after becoming overwhelmed with copies.
-NEWLINE_CHAR NEWLINE_CHAR The Oxfam shop in Swansea has been receiving an average of one copy of the Dan Brown novel a week for months, leaving them with little room for any other books.
-NEWLINE_CHAR NEWLINE_CHAR Staff who are struggling to sell copies of the book have put a note up in the store saying they would rather donors hand in their vinyl instead.
-– The Da Vinci Code has sold so many copies—that would be at least 80 million—that it’s bound to turn up in book donation piles.
-But at one charity shop in the UK, it’s been donated so heavily that the shop has posted a sign propped up on a tower of Da Vinci Code copies that reads: Ÿou could give us another Da Vinci Code… but we would rather have your vinyl!T̈he manager of the Oxfam shop in Swansea tells the Telegraph that people are laughing and taking pictures of the sizable display: Ï would say that we get one copy of the book every day.Ḧe says people buy them öccasionally,b̈ut with vinyl sales up 25% in the past year, they’d rather take records.
-Dan Brown’s book isn’t the only one that shops like Oxfam struggle to re-sell.
-Last year, Oxfam was hit with a large and steady supply of Fifty Shades of Grey, and it similarly begged donors: P̈lease—no more.B̈ut Brown has a particular kind of staying power.
-The Da Vinci Code was published in 2003, and within six years Brown had booted John Grisham from the No.
-1 slot on the list of writers whose books were most often donated to Oxfam’s 700 shops, reported the Guardian at the time.
-The Independent in 2012 reported Brown’s best-seller was the most-donated book for the fourth year running.
-(See why Dan Brown took heat from the Philippines.)
-Multi-XScience Despite the huge amount of recent research efforts on entity resolution (matching) there has not yet been a comparative evaluation on the relative effectiveness and efficiency of alternate approaches.
+[^8] https://docs.vllm.ai We followed model examples on how to interact for the summarization task (Table˜5).
+Appendix C Evaluation Details
+Multi-dimensional analyses with Prometheus.
+We developed rubrics for Prometheus based on definitions set forth by prior work.
+We altered the definitions slightly for the multi-document use case.
+For all criteria, i.e., description for model feedback, we use the same score rubric except for Factuality and Coverage:
+  1. Score 1:
+    Very poor
+  2. Score 2:
+    Poor
+  3. Score 3:
+    Barely acceptable
+  4. Score 4:
+    Good
+  5. Score 5:
+    Very Good
+The prompt format is in Figure˜9 and the rubrics are in Figure˜10.
+The rubric for Coverage, inspired by Wolhandler et al. (2022b):
+  1. Score 1:
+    The summary misses most salient information and/or fails to draw from many or all of the source documents.
+  2. Score 2:
+    The summary includes some salient information, but frequently omits key points or fails to adequately represent all source documents.
+  3. Score 3:
+    The summary covers a good portion of the salient information and represents most source documents, though some important details or content from some documents may be missing or underrepresented.
+  4. Score 4:
+    The summary incorporates most salient information from all source documents, with only minor omissions of less critical details, and adequately represents all documents.
+  5. Score 5:
+    The summary comprehensively and accurately captures all salient information from all source documents, providing a well-rounded representation that integrates content from all sources effectively.
+The rubric for Factuality, taken from the Prometheus code repository:
+  1. Score 1:
+    The model’s responses are mostly incorrect or based on unfounded information.
+  2. Score 2:
+    The model sometimes provides factually correct responses, but inaccuracies are common.
+  3. Score 3:
+    The model generally provides factually correct information, though some errors occur.
+  4. Score 4:
+    The model often provides factually accurate information with only occasional minor errors.
+  5. Score 5:
+    The model consistently provides responses that are factually correct and well-supported by evidence.
+Appendix D MDS Domain Generalizability
+We compare the results from training and testing on a single domain to training on a domain and testing on the others (“From”), and training on other domains and testing on the domain (“To”).
+“Ref” refers to the results when running the metrics on the ground-truth data.
+See Table˜3.
+Models trained on ConvoSumm-Reddit transfer well to other domains.
+On average, models trained on ConvoSumm-Reddit and tested on the News and Science datasets see minimal changes in performance across quality and ground-truth similarity metrics.
+The exception to this trend is the large decrease in MoverScore, which is more sensitive to a performance drop of a hundredth of a decimal due to its often low scores (e.g., 0.050.050.05).
+[^9] We discuss MoverScore further in Section 4.2.
+Transferring from Multi-News+ leads to a decrease in factuality.
+Compared to transferring from the other datasets, out-of-domain summaries generated with a model trained on Multi-News+ contain more factual errors (8-11% more).
+This is most likely due to the large style bias of smaller models, which tend to hallucinate quotes (see Figure˜13 for an example).
+This is different from the findings of (Belém et al., 2025), who found that models (LLMs) hallucinate more on conversation data than news data.
+Most metrics are impacted by models’ bias to the style of the training dataset.
+[…]
+Appendix E Raw Results
+The domain-transfer results in Table˜1 show the summarized relative performance of the models. The detailed performance across domains and models are in Table˜4.
+Example from Multi-News+
+Source   Documents
+Washington ( ap ) — as a crucial second sign-up season gears up, the obama administration said sunday that healthcare.gov is stable and working well, a far cry from last year’s frozen computer screens and frustrated customers.
+Health and human services secretary sylvia burwell smiles while answering a reporters question following her tour of the greater prince william community health center, evergreen terrace site, in manassas,...
+( associated press ) health and human services secretary sylvia burwell gestures while answering a reporters question following her tour of the greater prince william community health center, evergreen terrace site, in manassas,...
+( associated press ) health and human services secretary sylvia m.
+Burwell said she expects " strong and healthy growth " for 2015.
+About 7 million people are signed up, and burwell expects to grow that by 2 million more or so.
+The congressional budget office has projected a total of 13 million enrolled for 2015, and some see the administration as trying to lower expectations.
+Burwell told nbc’s " meet the press " that 100,000 people had submitted new applications this weekend via the federal website serving 37 states.
+That’s a big difference from last year, when only a handful of customers managed to enroll on the first day.
+Burwell also said that a half-million people who already have coverage through the program were able to log into their accounts this time.
+There were reports saturday that returning customers had problems, but some of that may have been confusion trying to remember user names and passwords.
+Administration spokesman aaron albright said sunday he had not seen any indication of problems.
+Burwell said call centers have taken 100,000 calls, another indication of consumer interest.
+President barack obama noted the improvements.
+" healthcare.gov works really well now, " he said.
+Healthcare.gov is an online marketplace that offers subsidized private coverage to people who don’t have health insurance on the job.
+Because of political opposition and
+Getty burwell: 100,000 new obamacare applications secretary of health and human services sylvia mathews burwell said on sunday 100,000 people submitted new applications for obamacare in the first days of the second open enrollment period.
+Speaking on nbc’s "meet the press, " burwell offered an update to the numbers provided saturday when she said 23,000 applied in the first eight hours of the new open enrollment period.
+Story continued below in addition to new applicants, 500,000 customers were able to log on to healthcare.gov, and 1 million people have gone "window shopping " to compare insurance plans and prices over the past week.
+The numbers offer a stark contrast to the first enrollment period a year ago when healthcare.gov launched with severe technical problems that prevented people from applying.
+Authors:
+Summary After last year’s dismal debut healthcare.gov is feeling much better now, and is busily connecting americans with health care plans as the second sign-up period opened yesterday, health and human services secretary sylvia burwell said today.
+The site took in 100,000 new applications in its first 24 hours, reports politico, saw 500,000 log in to their accounts, and 1 million more went " window-shopping " for plans.
+Call centers took 100,000 calls.
+Burwell said she expects " strong and healthy growth " this year — about another 2 million people to add to the 7 million already signed up, per the ap.
+A white house spokesman said today there was no indication of problems.
+" healthcare.gov works really well now, " said president obama.
+Figure 4: Example from the Multi-News+ dataset.
+Example from Multi-XScience
+Source Documents
+Despite the huge amount of recent research efforts on entity resolution (matching) there has not yet been a comparative evaluation on the relative effectiveness and efficiency of alternate approaches.
 We therefore present such an evaluation of existing implementations on challenging real-world match tasks.
@@
 It aims to identify similarities and differences across the techniques as well as their merits and limitations.
-For accurate entity linking and deduplication, early works on the subject examined many methods such as cosine similarity match, distance based match, TF IDF, and Soundex.
-The well known similarity measures for entity linking are summarized and reviewed in cite ; and the work in cite presents a comparative evaluation of some existing works.
-ConvoSumm Title: Hand from tonight.
+Summary For accurate entity linking and deduplication, early works on the subject examined many methods such as cosine similarity match, distance based match, TF IDF, and Soundex.
+The well known similarity measures for entity linking are summarized and reviewed in
+citekey; and the work in
+citekey presents a comparative evaluation of some existing works.
+Figure 5: Example from the Multi-XScience dataset.
+Example from ConvoSumm-Reddit
+Source Documents
+Title: Hand from tonight.
 Subreddit: poker Playing at my local pub league.
@@
 Sorry if this is bad formatting just quickly whipped it up for your opinions.
-|NEWLINE| Button: ( 30000) |NEWLINE| SB: Villian: (50000) …
+Button: ( 30000) SB: Villian: (50000) …
 Shove pre but I imagine this is the sort of game where you get told off for doing that
-A lot of commenters debate the semantics and word choice of another comment.
+Summary A lot of commenters debate the semantics and word choice of another comment.
 Some commenters say that the poster should shove even though other players will not like that move very much.
 Other commenters discuss the differences of some poker terminology.
-Table 3: Examples from each of the three datasets: Multi-News, Multi-XScience, and ConvoSumm-Reddit.
+Figure 6: Example from the ConvoSumm-Reddit dataset.
 # Examples # Source Docs. Summary Length Avg. Doc. Length
@@
 Val 50 16.74 (7.21) 59.18 (14.99) 52.88 (27.29)
-MultiNews Test 5621 2.73 (1.08) 217.00 (67.20) 982.98 (1779.55)
-Train 44964 2.76 (1.12) 218.27 (68.04) 982.95 (1851.20)
-Val 5621 2.70 (1.04) 216.72 (67.67) 980.06 (1163.94)
+MultiNews+ Test 4648 2.53 (0.88) 230.94 (69.57) 226.07 (64.15)
+Train 37057 2.55 (0.91) 233.54 (69.98) 224.30 (64.96)
+Val 4576 2.51 (0.86) 231.45 (69.64) 226.59 (63.35)
 MultiXScience Test 1901 3.55 (1.95) 106.35 (41.60) 195.32 (48.01)
@@
 Val 1937 3.55 (1.99) 104.03 (41.38) 196.98 (48.84)
-Table 4: Statistics for each dataset used in this work: Multi-News, Multi-XScience, and ConvoSumm-Reddit.
+Table 2: Statistics for each dataset used in this work: ConvoSumm-Reddit, Multi-News+, and Multi-XScience.
 The number of tokens in the (ground truth) summary and documents are measured with the BART tokenizer.
-Since we remove examples with missing information, our numbers might not match those in the original dataset papers (Appendix A).
-Domain Factuality Accuracy Quality
-Train/Test Cons. (UE) Rel. (UE) RL-Sum MoverScore Fluency (UE) Coh. (UE)
-ConvoSumm Convo 0.64 0.70 0.21 0.06 0.91 0.60
-News 0.64 0.62 0.19 0.17 0.89 0.66
-Science 0.67 0.55 0.15 0.01 0.69 0.46
-All 0.65 0.62 0.18 0.08 0.83 0.57
-PRIMERA Convo 0.65 0.73 0.21 0.09 0.92 0.71
-News 0.74 0.82 0.25 0.21 0.93 0.78
-Science 0.96 0.68 0.14 0.05 0.83 0.86
-All 0.78 0.74 0.20 0.12 0.89 0.78
-SummN Convo 0.65 0.63 0.21 0.07 0.89 0.62
-News 0.69 0.78 0.21 0.20 0.89 0.80
-Science 0.65 0.39 0.15 -0.02 0.66 0.27
-All 0.66 0.60 0.19 0.08 0.81 0.56
-Llama 3.1 Convo 0.80 0.92 0.14 0.07 0.93 0.94
-News 0.82 0.90 0.19 0.19 0.94 0.93
-Science 0.85 0.90 0.12 0.08 0.94 0.95
-All 0.83 0.91 0.13 0.08 0.93 0.95
-Table 5: Multi-document summarization (MDS) model performance when trained and tested on the same domain.
-The results for “All” are averaged across all domains for each model.
-Results are on the test splits of each dataset.
-Metrics followed by (UE) are from UniEval.
-For all metrics a greater score is better.
-Domain Factuality Accuracy Quality
-Train Test Fact. (UE) Rel. (UE) R1 RLSum MoverScore Fluency (UE) Coh. (UE)
-ConvoSumm Convo News 0.05 -0.41 -0.37 -0.44 -1.67 0.00 -0.25
-Convo Science 0.06 -0.09 -0.25 -0.33 -0.72 -0.01 0.02
-News Convo -0.11 0.30 -0.46 -0.31 -1.12 0.02 0.19
-News Science -0.07 0.10 -0.41 -0.28 -0.90 0.05 0.02
-Science Convo -0.07 -0.30 -0.16 -0.10 -2.79 0.03 -0.13
-Science News -0.01 -0.28 -0.01 -0.08 -1.90 0.07 -0.07
-All 0.01 -0.02 -0.11 -0.28 -0.26 -1.52 0.03 -0.04
-PRIMERA Convo News 0.20 -0.31 -0.05 -0.22 -0.22 -0.09 -0.08
-Convo Science 0.48 -0.07 -0.24 -0.36 -0.47 -0.10 0.20
-News Convo -0.22 0.00 -0.61 -0.57 -1.03 -0.07 0.00
-News Science 0.29 -0.17 -0.48 -0.45 -0.78 -0.11 0.09
-Science Convo -0.47 -0.12 -0.20 -0.13 -0.74 -0.27 -0.25
-Science News -0.11 0.04 0.43 0.30 2.25 -0.21 -0.03
-All 0.03 -0.10 -0.19 -0.24 -0.16 -0.14 -0.01
-SummN Convo News 0.31 -0.08 -0.22 -0.30 -0.67 0.04 0.14
-Convo Science 0.26 -0.02 -0.31 -0.35 -0.79 0.04 0.31
-News Convo -0.32 0.05 -0.69 -0.60 -1.22 -0.00 -0.00
-News Science -0.17 -0.21 -0.52 -0.41 -1.00 -0.01 -0.25
-Science Convo -0.00 -0.34 -0.59 -0.51 5.56 0.05 0.14
-Science News 0.17 0.03 0.37 0.17 -3.74 0.26 0.72
-All -0.06 0.04 -0.09 -0.32 -0.33 -0.31 0.07 0.18
-Llama 3.1 Convo News 0.04 -0.04 0.69 0.42 1.74 0.01 -0.01
-Convo Science 0.07 -0.03 0.05 0.03 0.29 -0.00 0.03
-News Convo -0.02 0.04 -0.48 -0.36 -0.72 -0.01 0.03
-News Science 0.05 -0.00 -0.42 -0.30 -0.56 -0.01 0.04
-Science Convo -0.06 0.02 -0.10 -0.05 -0.34 -0.01 -0.02
-Science News -0.03 -0.05 0.70 0.49 1.11 0.01 -0.05
-All 0.01 -0.01 0.07 0.04 0.25 -0.00 0.00
-Llama 3.1
-(0-shot)
-zero-shot Convo -0.01 -0.03 -0.16 -0.07 -0.23 -0.01 -0.02
-zero-shot News 0.02 -0.07 -0.01 0.06 -0.01 -0.00 -0.03
-zero-shot Science 0.01 -0.01 0.01 0.02 0.01 -0.01 0.01
-All 0.01 -0.04 -0.05 0.00 -0.07 -0.01 -0.01
-Table 6: Multi-document summarization (MDS) model relative performance when trained and tested on different domains.
-Results are on the test splits of each dataset.
-A positive number shows a performance gain and a negative number is a performance loss as compared to the model’s scores on the in-domain data.
-A relative change of over $1$ is possible with MoverScore because of its $[-1,1]$ range.
-Domain Factuality Accuracy Quality
-Train Test Fact. (UE) Rel. (UE) R1 RLSum MoverScore Fluency (UE) Coh. (UE)
-Convo News 0.15 -0.21 0.01 -0.14 -0.20 -0.01 -0.05
-Science 0.22 -0.05 -0.19 -0.25 -0.42 -0.02 0.14
-All 0.18 -0.13 -0.09 -0.19 -0.31 -0.01 0.04
-News Convo -0.17 0.10 -0.56 -0.46 -1.02 -0.02 0.05
-Science 0.03 -0.07 -0.46 -0.36 -0.81 -0.02 -0.03
-All -0.07 0.02 -0.51 -0.41 -0.92 -0.02 0.01
-Science Convo -0.15 -0.18 -0.26 -0.20 0.42 -0.05 -0.06
-News 0.01 -0.06 0.37 0.22 -0.57 0.03 0.14
-All -0.07 -0.12 0.06 0.01 -0.07 -0.01 0.04
-zero-shot Convo -0.01 -0.03 -0.16 -0.07 -0.23 -0.01 -0.02
-News 0.02 -0.07 -0.01 0.06 -0.01 -0.00 -0.03
-Science 0.01 -0.01 0.01 0.02 0.01 -0.01 0.01
-All 0.01 -0.04 -0.05 0.00 -0.07 -0.01 -0.01
-Table 7: Multi-document summarization (MDS) model relative performance when trained and tested on different domains.
-The results are averaged across all model types.
-Results are on the test splits of each dataset.
-Domain Factuality Accuracy Quality
-Train Test Cons. (UE) Rel. (UE) R1 RLSum MoverScore Fluency (UE) Coh. (UE)
-ConvoSumm Convo News 0.68 0.41 0.20 0.12 -0.04 0.92 0.45
-Convo Science 0.68 0.63 0.23 0.14 0.02 0.91 0.61
-News Convo 0.57 0.81 0.21 0.13 -0.02 0.90 0.78
-News Science 0.60 0.69 0.24 0.13 0.02 0.93 0.67
-Science Convo 0.63 0.38 0.21 0.14 -0.02 0.71 0.40
-Science News 0.67 0.39 0.25 0.14 -0.01 0.73 0.43
-PRIMERA Convo News 0.78 0.50 0.31 0.16 0.07 0.84 0.65
-Convo Science 0.96 0.68 0.25 0.14 0.05 0.83 0.86
-News Convo 0.58 0.82 0.18 0.11 -0.01 0.86 0.79
-News Science 0.96 0.68 0.25 0.14 0.05 0.83 0.86
-Science Convo 0.51 0.60 0.20 0.12 0.01 0.60 0.64
-Science News 0.86 0.71 0.35 0.18 0.15 0.65 0.83
-SummN Convo News 0.85 0.58 0.24 0.15 0.02 0.93 0.71
-Convo Science 0.82 0.62 0.22 0.14 0.01 0.92 0.81
-News Convo 0.47 0.82 0.13 0.08 -0.05 0.89 0.80
-News Science 0.57 0.62 0.21 0.12 -0.00 0.89 0.60
-Science Convo 0.64 0.26 0.10 0.07 -0.16 0.69 0.31
-Science News 0.76 0.40 0.33 0.18 0.07 0.83 0.47
-Llama 3.1 Convo News 0.83 0.89 0.41 0.22 0.18 0.94 0.93
-Convo Science 0.85 0.90 0.26 0.16 0.09 0.93 0.97
-News Convo 0.80 0.94 0.21 0.14 0.05 0.93 0.95
-News Science 0.86 0.90 0.24 0.15 0.08 0.94 0.96
-Science Convo 0.80 0.93 0.21 0.14 0.05 0.93 0.93
-Science News 0.82 0.86 0.39 0.23 0.18 0.94 0.91
-Llama 3.1 (0-shot) zero-shot Convo 0.79 0.90 0.21 0.15 0.05 0.93 0.92
-zero-shot News 0.83 0.84 0.40 0.23 0.19 0.94 0.90
-zero-shot Science 0.86 0.90 0.23 0.16 0.08 0.92 0.96
+Since we remove examples with missing information, our numbers might not match those in the original dataset papers (Appendix˜A).
+(a) ConvoSumm-Reddit
+(b) Multi-News+
+(c) Multi-XScience
+Figure 7: Pearson correlation between metrics on the test ground-truth (reference) for the evaluated datasets.
+(a) ConvoSumm
+(b) SummN
+(c) Llama 3.1 8B
+(d) PRIMERA
+Figure 8: Pearson correlation between metrics across all datasets for each model in the domain-transfer setting.
+ConvoSumm-Reddit Multi-News+ Multi-XScience
+From $\Delta$(%) To $\Delta$(%) Ref. From $\Delta$(%) To $\Delta$(%) Ref. From $\Delta$(%) To $\Delta$(%) Ref.
+Factuality Cons. (UE) 0.74 +11% -9% 0.71 0.80 -5% +8% 0.70 0.72 -5% +1% 0.55
+Fact. (P) 2.25 +11% -21% 3.07 2.46 -20% +10% 2.84 2.11 -5% -1% 1.82
+Cons. (P) 2.42 +5% -18% 3.06 2.62 -21% +11% 3.16 2.09 +1% -7% 1.86
+Ground-Truth Similarity Rel. (UE) 0.69 -1% -15% 0.99 0.73 0% -17% 0.98 0.63 -29% +5% 0.99
+R1 0.26 +13% -30% -1.00 0.42 -46% +28% -1.00 0.26 +6% -21% -1.00
+RLSum 0.17 -1% -22% -1.00 0.21 -37% +11% -1.00 0.15 +3% -20% -1.00
+MoverScore 0.05 -238% -98% 1.00 0.17 -80% -110% 1.00 0.05 -36% -107% 1.00
+Quality Rel. (P) 2.62 +5% -18% 3.23 2.81 -15% +8% 3.39 2.37 -4% -2% 2.78
+Fluency (UE) 0.85 +2% -5% 0.90 0.76 +4% -4% 0.76 0.77 -9% +7% 0.66
+Fluency (P) 2.69 +7% -16% 3.30 2.79 -21% +11% 2.75 2.17 +2% -5% 1.53
+Coh. (UE) 0.73 +6% -7% 0.73 0.81 -3% -3% 0.75 0.62 -13% +2% 0.43
+Coh. (P) 2.60 +6% -19% 3.39 2.89 -24% +12% 3.70 2.17 +2% -7% 1.51
+Comp.-Sent 9.01 +45% -72% 12.03 2.74 -44% +19% 2.08 4.07 -42% +18% 6.10
+Comp.-Word 7.18 +13% -24% 8.63 2.47 -18% +8% 1.98 4.40 -8% +4% 5.92
+Table 3: Multi-document summarization (MDS) model performance and relative performance aggregated by domain when trained and tested on the same domain, and the relative change ($\Delta$) when transferred to other domains (From), and vice-versa (To).
+A relative change of over 111 is possible with MoverScore because of its [1,1][-1,1][ - 1 , 1 ] range.
+For all metrics other than Compression, a higher score is better.
+Metrics measured by UniEval and Prometheus are marked by (UE) and (P), respectively.
+Domain Factuality Ground-Truth Similarity Quality
+Train Test C(UE) F(P) C(P) R(UE) R1 RLS MS R(P) Fl(UE) Fl(P) Co(UE) Co(P) Cov(P) C.Sent C.Word
+ConvoSumm Convo News+ 0.75 1.83 1.90 0.51 0.24 0.14 0.00 2.08 0.91 2.02 0.57 1.94 1.68 -4.31 -7.54
+Convo Science 0.67 1.95 1.97 0.61 0.23 0.14 0.02 2.27 0.91 2.22 0.58 2.03 1.76 -5.17 -9.23
+News+ Convo 0.68 1.59 1.78 0.75 0.23 0.13 0.01 1.97 0.70 1.72 0.78 1.74 1.77 -3.62 -2.51
+News+ Science 0.69 1.77 1.76 0.70 0.23 0.13 0.04 2.13 0.81 1.79 0.71 1.87 1.89 -2.31 -2.32
+Science Convo 0.44 1.22 1.26 0.26 0.22 0.15 -0.01 1.34 0.60 1.27 0.26 1.29 1.25 -9.32 -7.78
+Science News+ 0.66 1.81 1.87 0.26 0.30 0.16 0.04 2.10 0.64 1.65 0.30 1.70 1.69 -3.37 -3.99
+PRIMERA Convo News+ 0.83 1.97 2.08 0.41 0.38 0.18 0.13 2.38 0.63 2.04 0.71 1.94 2.01 -1.99 -1.28
+Convo Science 0.84 1.96 1.95 0.61 0.23 0.12 0.06 2.38 0.79 1.82 0.73 1.79 1.94 -2.23 -1.59
+News+ Convo 0.82 1.34 1.53 0.40 0.19 0.11 -0.01 1.76 0.67 1.36 0.72 1.26 1.52 -5.29 -3.17
+News+ Science 0.82 2.03 1.96 0.65 0.23 0.12 0.06 2.40 0.82 1.92 0.74 1.83 2.04 -2.40 -1.61
+Science Convo 0.80 1.43 1.49 0.41 0.19 0.11 -0.02 1.77 0.64 1.34 0.67 1.28 1.58 -4.46 -2.61
+Science News+ 0.78 1.96 2.07 0.44 0.38 0.18 0.13 2.33 0.62 1.95 0.68 1.94 1.99 -2.42 -1.51
+SummN Convo News+ 0.90 2.71 2.64 0.64 0.23 0.14 0.01 2.75 0.89 2.92 0.82 2.74 2.18 -5.47 -8.28
+Convo Science 0.77 2.51 2.36 0.72 0.21 0.13 0.01 2.48 0.92 2.61 0.88 2.27 1.99 -7.72 -11.28
+News+ Convo 0.67 1.44 1.58 0.71 0.22 0.13 0.00 1.95 0.70 1.74 0.73 1.58 1.67 -3.45 -2.16
+News+ Science 0.64 1.57 1.54 0.67 0.21 0.12 0.03 2.15 0.77 1.63 0.68 1.68 1.83 -1.39 -1.26
+Science Convo 0.39 1.13 1.21 0.33 0.21 0.14 -0.04 1.26 0.63 1.37 0.33 1.25 1.26 -6.52 -4.14
+Science News+ 0.65 1.95 2.04 0.33 0.31 0.17 0.04 2.36 0.69 2.05 0.35 2.07 1.90 -3.14 -3.58
+Llama 3.1 Convo News+ 0.87 3.24 3.56 0.92 0.40 0.21 0.16 3.64 0.94 4.11 0.95 3.98 3.02 -2.31 -1.79
+Convo Science 0.88 3.32 3.47 0.91 0.24 0.15 0.08 3.58 0.94 4.06 0.97 4.00 3.23 -1.39 -1.29
+News+ Convo 0.82 3.01 3.23 0.94 0.22 0.15 0.05 3.43 0.94 3.99 0.96 3.92 2.96 -3.88 -1.36
+News+ Science 0.88 3.38 3.45 0.92 0.24 0.15 0.09 3.61 0.93 4.03 0.98 4.00 3.19 -1.30 -1.26
+Science Convo 0.82 3.17 3.21 0.94 0.22 0.15 0.06 3.47 0.93 3.94 0.96 3.92 2.91 -4.00 -1.46
+Science News+ 0.87 3.25 3.54 0.94 0.40 0.22 0.16 3.64 0.95 4.07 0.95 3.98 2.93 -2.15 -1.69
+Llama 3.1 (0-shot) zero-shot Convo 0.83 3.12 3.31 0.94 0.22 0.15 0.05 3.44 0.93 4.01 0.96 3.87 2.79 -3.74 -1.40
+zero-shot News+ 0.88 3.26 3.41 0.93 0.41 0.22 0.16 3.57 0.94 4.04 0.94 4.00 3.02 -2.07 -1.64
+zero-shot Science 0.87 3.42 3.52 0.93 0.24 0.16 0.08 3.47 0.93 4.06 0.98 3.96 3.29 -1.38 -1.30
 Table 4: Multi-document summarization (MDS) model performance when trained and tested on the different domains.
 Results are on the test splits of each dataset.
-Metrics followed by (UE) are from UniEval.
-For all metrics a greater score is better.
-Stage Setting Multi-News ConvoSumm-Reddit Multi-XScience
+For all metrics, a greater score is better.
+Prometheus Evaluator Prompt
+System You are a fair judge assistant tasked with providing clear, objective feedback based on specific criteria, ensuring each assessment reflects the absolute standards set for performance.
+Assistant ###Task Description: An instruction (might include an Input inside it), a response to evaluate, a reference answer that gets a score of 5, and a score rubric representing a evaluation criteria are given.
+1.
+Write a detailed feedback that assess the quality of the response strictly based on the given score rubric, not evaluating in general.
+2.
+After writing a feedback, write a score that is an integer between 1 and 5.
+You should refer to the score rubric.
+3.
+The output format should look as follows: "(write a feedback for criteria) [RESULT] (an integer number between 1 and 5)" 4.
+Please do not generate any other opening, closing, and explanations.
+###The instruction to evaluate: You are a multi-document summarizer.
+When given a list of DOCUMENTS, you provide a SUMMARY that incorporates all of the documents.
+Write a summary that incorporates all of the following documents: ### DOCUMENTS {all source documents} ###Response to evaluate: {model-generated summary} ###Reference Answer (Score 5): {ground-truth summary} ###Score Rubrics: {rubric} ###Feedback:
+Figure 9: The formatting is from Kim et al. (2024), and the instructions and rubric are from Figure˜10.
+Evaluation Criteria for Prometheus
+Factuality Is the model’s summary factually correct and well-supported by the DOCUMENTS?
+Relevance How well does the summary capture the key points of the documents?
+Consider whether all and only the important aspects are contained in the summary.
+Only evaluate the summary for relevance.
+Do not penalize the summary for issues that are outside the scope of relevance.
+Consistency How consistent are the facts in the summary with the facts in the original documents?
+Consider whether the summary does reproduce all facts accurately and does not make up untrue information.
+Do not use your own knowledge, and consider the documents as truth.
+Only evaluate the summary for consistency.
+Do not penalize the summary for issues that are outside the scope of consistency.
+Fluency What is the quality of individual sentences, are they well-written and grammatically correct?
+Consider the quality of individual sentences.Only evaluate the summary for fluency.
+Do not penalize the summary for issues that are outside the scope of fluency.
+Coherence What is the quality of all sentences collectively?
+How well do they fit together, and do they sound natural?
+Consider the quality of the summary as a whole.
+Only evaluate the summary for coherence.
+Do not penalize the summary for issues that are outside the scope of coherence.
+Coverage Does the model’s summary incorporate salient information from all the documents? Do not penalize the summary for issues that are outside the scope of coverage.
+Figure 10: Evaluation dimensions provided to Prometheus (Kim et al., 2024).
+The criteria for Relevance, Consistency, Fluency, and Coherence are drawn from SummEval (Fabbri et al., 2021b).
+We added the italicized text to the prompt to improve Prometheus performance.
+Evaluation Criteria for UniEval (Summarization)
+Relevance Is this summary relevant to the reference?
+Consistency Is this claim consistent with the document?
+Fluency Is this a fluent paragraph?
+Coherence Is this a coherent summary to the document?
+Figure 11: Evaluation dimensions for UniEval (Zhong et al., 2022). Only Relevance requires access to a reference summary.
+Llama 3.1 Instruct Prompt Templates
+Zero-shot
+<|begin_of_text|><|start_header_id|>user<|end_header_id|>
+Write a summary that incorporates all of the following documents:
+Document: {source document 1}
+Document: {source document $n$}
+<|eot_id|><|start_header_id|>assistant<|end_header_id|>
+In-context Learning
+<|begin_of_text|><|start_header_id|>user<|end_header_id|>
+Write a summary that incorporates all of the following documents:
+Document: {source document 1}
+Document: {source document $n$}
+<|eot_id|><|start_header_id|>assistant<|end_header_id|>
+{target summary}
+<|begin_of_text|><|start_header_id|>user<|end_header_id|>
+Write a summary that incorporates all of the following documents:
+Document: {source document 1}
+Document: {source document $n$}
+<|eot_id|><|start_header_id|>assistant<|end_header_id|>
+Table 5: Prompt templates for Llama 3.1 8B Instruct for the zero-shot (top) and single-shot in-context learning tasks (bottom).
+Stage Setting Multi-News+ ConvoSumm-Reddit Multi-XScience
 General
@@
 beam 1 - -
-length penalty 2 1 -
+length penalty 1 - -
 max length 450 - -
-min length 150 0 -
-no-repeat-ngram-size 3 0 -
+min length 0 - -
+no-repeat-ngram-size 3 - -
+temperature 0.9 - -
 Stage 2
@@
 beam 1 - -
-length penalty 2 1 -
+length penalty 1 - -
 max length 550 - -
-min length 150 0 -
-no-repeat-ngram-size 3 0 -
+min length 0 - -
+no-repeat-ngram-size 3 - -
+temperature 0.9 - -
 Stage 3
@@
 beam 1 - -
-length penalty 2 1 -
+length penalty 1 - -
 max length 900 1024 -
-min length 400 0 -
-no-repeat-ngram-size 3 0 -
+min length 0 - -
+no-repeat-ngram-size 3 - -
+temperature 0.9 - -
 Table 6: Training details for SummN on all datasets. Weights initialized from BART-Large-CNN. An entry of “-” indicates the same setting as the previous column.
-Setting Multi-News* ConvoSumm-Reddit Multi-XScience*
+Setting Multi-News+ ConvoSumm-Reddit Multi-XScience*
 max input sequence 4096 4096 4096
-total number of steps 25K 5K 20K
-warmup steps 2.5K 200 2K
+total number of steps 30K 5K 20K
+warmup steps 1K 200 2K
 optimizer Adam Adam Adam
@@
 Table 7: Training details for PRIMERA on all datasets.
-Models marked with (*) were trained by (Xiao et al., 2022b) and only available training information is filled in.
+Models marked with (*) were trained by (Xiao et al., 2022) and only available training information is filled in.
 Weights initialized from BART-Large.
-Setting Multi-News ConvoSumm-Reddit* Multi-XScience
+Setting Multi-News+ ConvoSumm-Reddit* Multi-XScience
 max input sequence 2048 2048 2048
@@
 An entry of “-” indicates the same setting as the previous column.
-Zero-shot
-<|begin_of_text|><|start_header_id|>user<|end_header_id|>
-Write a summary that incorporates all of the following documents:
-Document: {source document 1}
-Document: {source document $n$}
-<|eot_id|><|start_header_id|>assistant<|end_header_id|>
-In-context Learning
-<|begin_of_text|><|start_header_id|>user<|end_header_id|>
-Write a summary that incorporates all of the following documents:
-Document: {source document 1}
-Document: {source document $n$}
-<|eot_id|><|start_header_id|>assistant<|end_header_id|>
-{target summary}
-<|begin_of_text|><|start_header_id|>user<|end_header_id|>
-Write a summary that incorporates all of the following documents:
-Document: {source document 1}
-Document: {source document $n$}
-<|eot_id|><|start_header_id|>assistant<|end_header_id|>
-Table 12: Prompt templates for Llama 3.1 8B Instruct for the zero-shot (top) and single-shot in-context learning tasks (bottom).
+Generated Summaries: PRIMERA Models
+Multi-XScience Comp-Word: 3.06
+R-LSum: 0.16
+Fact. (P): 2.00
+Fact. (U): 0.80
+Fluency (P): 2.00
+Fluency (U): 0.53
+Washington ( ap ) — as a crucial second sign-up season gears up, the obama administration said sunday that healthcare.gov is stable and working well, a far cry from last year’s frozen computer screens and frustrated customers.
+Health and human services secretary sylvia burwell smiles while answering a reporters question following her tour of the greater prince william community health center, evergreen terrace site, in manassas,...
+( associated press ) health and human service secretary syLVia burrell gestures while answeringa reporters question after her tour...
+( associated Press ) health care secretary sylvia burwwind is an employee of the federal government who works to make healthcare....
+(associated press ) Getty.Getty burwell: 100,000 new obamacare applications secretary of health and h...
+(Associated press ) healthcare.hgov is an online marketplace that offers subsidized private coverage to people who don’t have health insurance on the job.Because of political opposition and the
+ConvoSumm Comp-Word: 2.04
+R-LSum: 0.18
+Fact. (P): 2.00
+Fact. (U): 0.91
+Fluency (P): 2.00
+Fluency (U): 0.67
+Washington ( ap ) — as a crucial second sign-up season gears up, the obama administration said sunday that healthcare.gov is stable and working well, a far cry from last year’s frozen computer screens and frustrated customers.
+Health and human services secretary sylvia burwell smiles while answering a reporters question following her tour of the greater prince william community health center, evergreen terrace site, in manassas,...
+( associated press ) health and human rights secretary sylivia burforth gestures while answering her tour in the greater Prince william Community health center during the first days of the second open enrollment period.Health and Human Services Secretary Sylvia Mathews Burwell also said a half-million people who already have coverage through the program were able to log into their accounts this time, compared to last year when thousands struggled to do so.Burwell said call centers have taken 100,000 calls, another indication of consumer interest.President barack obama noted the improvements.Healthcare.gov’s numbers were down slightly from last season’s first open enrollment.Officials said last year many people would not sign up because of the website problems.Aaron Albright, an Obama administration spokesman, said health and Human services’ healthcare.org site " is accessible for anyone seeking to access the site" and the agency has been "redesigned and upgraded" for this year.Last week, the site’s status improved significantly.Reuters contributed to this report..
+MultiNews+ Comp-Word: 2.80
+R-LSum: 0.19
+Fact. (P): 3.00
+Fact. (U): 0.83
+Fluency (P): 3.00
+Fluency (U): 0.74
+Washington ( ap ) — as a crucial second sign-up season gears up, the obama administration said sunday that healthcare.gov is stable and working well, a far cry from last year’s frozen computer screens and frustrated customers.
+Health and human services secretary sylvia burwell smiles while answering a reporters question following her tour of the greater prince william community health center, evergreen terrace site, in manassas,...
+( associated press ) health and humanservices secretary syvla m.
+Burwell gestures while answering the reporters question after her tour on the greater Prince william Community Health Center, Evergreen Terrace site in Manassas, Va.
+( associated Press ) health care.gov website in Washington, D.C.
+( Associated Press ) In an interview with NBC, Burwell said: "Healthcare.gov works really well now," adding that "in addition to new applicants," 500,000 people were able to log on to healthcare.com this week, and 1 million people have gone "window shopping" to compare insurance plans and prices over the past week.
+Figure 12: Example generated summaries from PRIMERA models for a Multi-News+ example. Source documents and ground-truth reference summary are in Figure˜4.
+Generated Summaries: SummN Models
+MultiXScience Comp-Word: 6.86
+R-LSum: 0.20
+Fact. (P): 2.00
+Fact. (U): 0.93
+Fluency (P): 3.00
+Fluency (U): 0.80
+The secretary of health and human services reported that 100,000 people submitted new applications for Obamacare in the first days of the second open enrollment period @cite .
+The secretary also said that call centers have taken 100,00 calls, another indication of consumer interest.
+The administration has projected a total of 13 million enrolled for 2015, and some see the administration as trying to lower expectations.
+ConvoSumm Comp-Word: 6.04
+R-LSum: 0.21
+Fact. (P): 3.00
+Fact. (U): 0.97
+Fluency (P): 3.00
+Fluency (U): 0.96
+The secretary of health and human services said that 100,000 people submitted new applications for Obamacare in the first days of the second open enrollment period.
+She also said that a half-million people who already have coverage through the program were able to log into their accounts this time.
+The numbers offer a stark contrast to the first enrollment period a year ago when healthcare.gov launched with severe technical problems that prevented people from applying.
+MultiNews+ Comp-Word: 3.91
+R-LSum: 0.24
+Fact. (P): 2.00
+Fact. (U): 0.44
+Fluency (P): 3.00
+Fluency (U): 0.76
+The second open enrollment period for obamacare’s 2015 enrollment period kicked off today, and the administration says it’s already seeing a big jump in applications.
+Health and human services secretary sylvia burwell told meet the press that 100,000 people applied for coverage on the federal website this weekend, up from 23,000 on the first day last year, reports politico.
+" healthcare.gov works really well now, " said burwell.
+The administration expects " strong and healthy growth " for 2015, she said, adding that 7 million people have signed up so far, and she expects that number to grow by 2 million more.
+The congressional budget office has projected a total of 13 million enrolled for 2015.
+Figure 13: Example generated summaries from SummN models for a Multi-News+ example. Source documents and ground-truth reference summary are in Figure˜4.
+Generated Summaries: ConvoSumm Models
+MultiXScience Comp-Word: 8.55
+R-LSum: 0.19
+Fact. (P): 3.00
+Fact. (U): 0.21
+Fluency (P): 1.00
+Fluency (U): 0.62
+@cite @cite , the secretary of health and human services @cites said that 100,000 people submitted new applications for obamacare in the first days of the second open enrollment period.
+@ccite also said that a half-million people who already have coverage through the program were able to log into their accounts this time.
+ConvoSumm-Reddit Comp-Word: 14.61
+R-LSum: 0.08
+Fact. (P): 1.00
+Fact. (U): 0.88
+Fluency (P): 2.00
+Fluency (U): 0.96
+Most commentators say that the website is stable and working well. One commentator says that the number of new applicants is a good sign. Another commentator says the website looks good.
+MultiNews+ Comp-Word: 2.44
+R-LSum: 0.20
+Fact. (P): 3.00
+Fact. (U): 0.92
+Fluency (P): 3.00
+Fluency (U): 0.82
+Health and human services secretary sylvia burwell says healthcare.gov is stable and working well, a far cry from last year’s frozen computer screens and frustrated customers.
+Burwell told nbc’s meet the press that 100,000 people had submitted new applications this weekend via the federal website serving 37 states.
+That’s a big difference from last week, when only a handful of customers managed to enroll on the first day.
+Bur Well also said that a half-million people who already have coverage through the program were able to log into their accounts this time.
+There were reports saturday that returning customers had problems, but some of that may have been confusion trying to remember user names and passwords.
+Administration spokesman aaron albright said sunday he had not seen any indication of problems.
+Bur well said she expects " strong and healthy growth " for 2015.
+About 7 million people are signed up, and burwell expects to grow that by 2 million more or so.
+The congressional budget office has projected a total of 13 million enrolled for 2015, and some see the administration as trying to lower expectations.
+Figure 14: Example generated summaries from ConvoSumm models for a Multi-News+ example. Source documents and ground-truth reference summary are in Figure˜4.
+Generated Summary: Llama 3.1 Zero-Shot
+Zero-Shot Comp-Word: 3.00
+R-LSum: 0.25
+Fact. (P): 4.00
+Fact. (U): 0.86
+Fluency (P): 4.00
+Fluency (U): 0.96
+As the second sign-up season for healthcare under the Obama administration gears up, officials report significant improvements over last year’s struggles.
+According to Health and Human Services Secretary Sylvia Burwell, approximately 7 million people are currently signed up, with an expectation to grow by 2 million more or so.
+This year’s enrollment progress is a stark contrast to last year’s frozen computer screens and frustrated customers.
+In the first days of the current enrollment period, 100,000 new applications have been submitted via the federal website, healthcare.gov, which serves 37 states.
+An additional 500,000 existing customers were able to access their accounts, and 1 million individuals have explored insurance plans and prices on the website.
+President Barack Obama has noted the improvement, stating that "healthcare.gov works really well now." The administration expects "strong and healthy growth" for 2015, with a projected total of 13 million enrolled, surpassing the previous year’s enrollment figures.
+Figure 15: Example generated summary from the Zero-Shot model for a Multi-News+ example. Source documents and ground-truth reference summary are in Figure˜4.
+Generated Summaries: In-Context Learning (ICL) with Llama 3.1 Models
+ConvoSumm-Reddit Comp-Word: 3.06
+R-LSum: 0.31
+Fact. (P): 4.00
+Fact. (U): 0.83
+Fluency (P): 5.00
+Fluency (U): 0.96
+As the second sign-up season for healthcare under the Obama administration gears up, officials are reporting significant improvements in the stability and functionality of healthcare.gov.
+Health and Human Services Secretary Sylvia Burwell has stated that the website is now working well, with over 100,000 new applications submitted in the first weekend, compared to just a handful of customers who managed to enroll last year.
+Additionally, approximately 500,000 customers were able to log into their existing accounts, and over 1 million people have begun "window shopping" to compare insurance plans and prices.
+These numbers are in stark contrast to the technical difficulties and frustrations experienced during the first enrollment period, and suggest a strong and healthy growth for the program in 2015, with the Congressional Budget Office projecting a total of 13 million enrolled.
+President Barack Obama has also noted the improvements, stating that healthcare.gov now works "really well."
+Multi-XScience Comp-Word: 2.45
+R-LSum: 0.27
+Fact. (P): 3.00
+Fact. (U): 0.90
+Fluency (P): 4.00
+Fluency (U): 0.96
+As the second sign-up season for healthcare under the Obama administration gears up, officials are reporting significant improvements in the stability and functionality of healthcare.gov.
+Health and Human Services Secretary Sylvia Burwell has stated that the website is now working well, in stark contrast to last year’s technical issues that plagued enrollment efforts.
+According to Burwell, over 100,000 new applications have been submitted via the federal website, serving 37 states, since the beginning of the second open enrollment period.
+This is a significant increase from last year, when only a handful of customers managed to enroll on the first day.
+Additionally, over 500,000 existing customers have been able to log into their accounts, and 1 million people have gone "window shopping" to compare insurance plans and prices over the past week.
+The numbers suggest a high level of consumer interest and a continued growth in enrollment.
+The Congressional Budget Office has projected a total of 13 million enrolled for 2015, and Burwell expects to reach this goal or come close.
+President Barack Obama has also noted the improvements, stating that "healthcare.gov works really well now."
+MultiNews+ Comp-Word: 2.49
+R-LSum: 0.25
+Fact. (P): 3.00
+Fact. (U): 0.94
+Fluency (P): 4.00
+Fluency (U): 0.96
+As the second sign-up season for the Affordable Care Act (ACA) kicks off, the Obama administration is reporting significant improvements in the healthcare.gov website.
+Health and Human Services Secretary Sylvia Burwell has stated that the website is stable and working well, in stark contrast to last year’s technical issues that frustrated customers.
+Despite some initial concerns about returning customers experiencing problems logging into their existing accounts, Burwell has reported that 500,000 people were able to access their accounts on the website.
+In the first weekend of the new open enrollment period, 100,000 new applications were submitted through healthcare.gov, with an additional 100,000 calls made to call centers.
+Additionally, 1 million people have gone "window shopping" to compare insurance plans and prices over the past week.
+These numbers represent a significant increase from last year’s enrollment period, where only a handful of customers managed to enroll on the first day.
+Burwell has expressed confidence in the website’s ability to handle the increased traffic, stating that she expects "strong and healthy growth" for 2015, with 13 million people projected to enroll through the website.
+Figure 16: Example generated summaries from In-Context Learning (ICL) models for a Multi-News+ example. Source documents and ground-truth reference summary are in Figure˜4.